### v1

# See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models

Abstract
[…]
See or Say Graphs: Agent-Driven Scalable Graph Understanding
with Vision-Language Models
Shuo Han1,2*   Yukun Cao1,2*   Zezhong Ding1,2 Zengyi Gao1,2S. Kevin Zhou1,3Xike Xie1,2 1University of Science and Technology of China, China 2Data Darkness Lab, MIRACLE Center, USTC, China 3MIRACLE Center, Suzhou Institute for Advance Research, USTC, China {shuo.han, ykcho, zezhongding, gzy02}@mail.ustc.edu.cn {skevinzhou, xkxie}@ustc.edu.cn
**footnotetext: These authors contributed equally to this work.footnotetext: Corresponding author.
1 Introduction
Recently, vision-language models (VLMs) have shown potential for general-purpose understanding of graph data Tang et al. (2024a); Kong et al. (2025b); Ding et al. (2025b), offering new paradigms for solving real-world problems naturally represented as graphs. A central research focus of this emerging field is to enhance the foundational capability of VLMs to understand graph structures Ren et al. (2024), particularly for large-scale graphs, without relying on external tools such as code execution or software systems Ding et al. (2025a); Feng et al. (2025). To this end, existing studies have explored two primary graph input modalities for VLMs: textual graph descriptions [^1] Many early studies on text-modality-based graph understanding were developed on LLMs. As VLMs extend LLMs with visual capabilities, this paper presents a unified discussion from the perspective of graph understanding with VLMs., which convert structural information into natural language for indirect memory and reasoning over node properties and relations Jiang et al. (2023); Jin et al. (2024); Zhang et al. (2024a); and visual graph representations, which render nodes and edges as structured images to support direct structural perception Li et al. (2024c); Zhu et al. (2025). More recent efforts further combine these modalities to improve structural understanding and complex reasoning on graphs Wei et al. (2024).
Figure 1: Comparison of textual and visual modalities across different types of graph understanding tasks
However, despite these advances, current VLMs still lack a unified framework that systematically addresses the scalability challenges and the difficulties in coordinating textual and visual modalities inherent in graph understanding. Specifically:
Scalability remains a fundamental limitation for both textual and visual graph inputs in VLMs. Given their limited input token capacities, textual graph descriptions are constrained by sequence length, making it difficult to encode the full structure of large graphs Chen et al. (2024a); Yuan et al. (2025a). Visual graph inputs, on the other hand, are restricted by image resolution, resulting in the loss of structural details as graph size increases Zhu et al. (2025).
In addition, VLMs face the lack of an effective mechanism to coordinate different input modalities for graph understanding, which prevents them from fully leveraging the distinct strengths and complementary advantages of textual and visual modalities. As illustrated in Figure 1, under the token length limits of VLMs, the text modality is more effective for simple property tasks such as counting nodes, measuring degrees, and assessing connectivity. In these cases, the information can be directly located and reasoned over with minimal effort. By contrast, the visual modality often performs worse because the limited resolution of global visual graphs blurs structural details, causing VLMs to misidentify relevant elements and leading to performance degradation.
For tasks that require local and structurally complex reasoning, such as shortest-path computation or cycle detection, high-resolution local visual subgraphs offer a clear advantage for VLMs. They provide a “what-you-see-is-what-you-get” view of the topology and naturally support stepwise reasoning. In contrast, textual descriptions are often lengthy and lack structural clarity, leaving VLMs with only indirect language cues and thus little structural grounding for effective complex graph reasoning.
Therefore, a natural conclusion is that realizing the full potential of VLMs in graph understanding requires addressing these two challenges.
This entails, on the one hand, preserving as much task-relevant graph information as possible within limited input tokens, and on the other hand, assigning tasks according to modality strengths: letting the text modality efficiently address simple property tasks, while leveraging the visual modality to directly support local and structurally complex tasks.
To this end, we propose GraphVista, a unified framework for graph understanding with VLMs. For the scalability challenges, GraphVista is inspired by the retrieval-augmented generation (RAG) paradigm Lewis et al. (2020). It organizes graph information hierarchically according to structural importance and stores multi-granularity textual descriptions in a lightweight GraphRAG base. For a given graph understanding task, GraphVista retrieves only the relevant information before input, thereby compressing redundant context and preserving key reasoning elements. When necessary, it can also generate high-resolution visual subgraphs from the retrieved local information to support complex reasoning tasks.
To enable effective coordination across modalities, GraphVista introduces a planning agent that parses graph understanding tasks and routes them to the appropriate modality. The textual modality branch, enhanced with the RAG techniques, efficiently handles simple property tasks by retrieving and reasoning over concise textual descriptions. In contrast, the visual modality branch employs the Visual Graph Thoughts Agent, which constructs a multimodal chain of thought called Visual Graph Thoughts over high-resolution visual subgraphs, enabling VLMs to perform stepwise visual reasoning grounded in explicit structural evidence for structurally complex tasks.
The main contributions of this paper are listed as follows:
  • $\bullet$
    We present the first systematic analysis of the scalability challenges and coordination gaps that hinder VLMs in graph understanding, providing both conceptual insights and empirical evidence.
  • $\bullet$
    We propose GraphVista, the first unified framework for graph understanding with VLMs, designed around two objectives: improving scalability and enabling coordination across modalities. To this end, GraphVista integrates a set of innovative techniques, such as a hierarchical GraphRAG base, a planning agent for task routing, and a Multimodal Graph Thoughts agent for complex reasoning.
  • $\bullet$
    We introduce Grena, a scalable benchmark that supports step-level training of Visual Graph Thoughts and evaluates graph understanding from simple property to complex local tasks.
  • $\bullet$
    Extensive experiments across diverse graph understanding tasks demonstrate the effectiveness of GraphVista, highlighting its ability to scale to large graphs and to exploit the complementary strengths of textual and visual modalities.
2 Related Work
2.1 Text Modality-based Graph Understanding
Most existing studies convert graphs into textual descriptions to enable VLM-based graph understanding, primarily focusing on benchmark construction and capability analysis. Recent benchmarks evaluate VLMs on basic graph element comprehension Wang et al. (2023); Guo et al. (2023); Wu et al. (2024) and graph-theoretic reasoning tasks Chen et al. (2024a); Li et al. (2024b); Tang et al. (2024b); Luo et al. (2024); Yuan et al. (2025a), indicating that the research in this direction remains in an early exploratory phase. Methodologically, Ge et al. (2025); Wang et al. (2025) studied the influence of graph description order, while Cao et al. (2025) identified the “Lost in the Middle” issue in graph sequence. Building upon these findings, Guan et al. (2025) analyzed VLM attention patterns over graph data, and Peng et al. (2025); Zhang et al. (2025) explored task transferability to broader reasoning domains. However, these strategies still face scalability bottlenecks and struggle with complex reasoning tasks.
Other studies address graph problems by leveraging external tools. Li et al. (2024a); Yuan et al. (2025b) employ predefined coding templates for algorithmic reasoning, while Perozzi et al. (2024) trains task-specific GNNs for graph processing. These methods depend on external modules rather than enhancing the intrinsic graph understanding of VLMs, and are thus orthogonal to our work.
2.2 Visual Modality-based Graph Understanding
Visual modality-based graph understanding is still in its infancy. VisionGraph Li et al. (2024c) and VGCure Zhu et al. (2025) benchmark VLMs on graph perception and reasoning, revealing difficulties in capturing structural information. Wei et al. (2024) further introduces a hybrid multimodal representation directly compatible with VLMs. However, these studies remain empirical and fail to address the scalability, structural understanding, and reasoning limitations inherent to vision-only representations.
3 Methodology
Figure 2: Framework of GraphVista
3.1 Overview
In this section, we introduce GraphVista, an agent-driven framework designed to address the challenges of scalability and modality coordination in graph understanding, as shown in Figure 2.
GraphVista operates under the coordination of the planning agent, which orchestrates the entire reasoning workflow. Initially, the planning agent constructs a lightweight Hierarchical GraphRAG Base $\mathcal{K}$ by partitioning the graph based on topological centrality. Upon receiving a question, it parses the task and routes it to the appropriate reasoning modality. For Simple Property Tasks, the text modality-based branch is activated, employing GraphRAG-based retrieval to extract relevant context from $\mathcal{K}$. For Complex Local Tasks that require fine-grained structural reasoning, the task is delegated to the visual modality-based branch. The Visual Graph Thoughts Agent extracts visual subgraphs from $\mathcal{K}$ and performs iterative multimodal reasoning grounded in visual evidence.
The rest of this section details our methodology. We begin by introducing the construction of $\mathcal{K}$ in §3.2. Next, we describe the planning agent-based task routing mechanism in §3.3. Finally, we elaborate on the visual modality-based and text modality-based graph understanding modules in §3.4 and §3.5, respectively.
3.2 Hierarchical GraphRAG Base
A primary challenge for LLMs and VLMs in understanding graph structures is their limited input capacity. This constraint makes it infeasible to encode an entire graph, often causing a loss of critical structural information. To address this limitation, we draw inspiration from RAG and propose building $\mathcal{K}$. Unlike conventional RAG pipelines Guo et al. (2025); Jin et al. (2025) that rely on external knowledge sources, our approach constructs a compact multi-level structural representation directly from the graph and stores it in a lightweight GraphRAG Base, enabling scalable reasoning without external dependencies [^2] For small graphs with $|V|\leq 15$, we skip storage and feed the entire graph directly into the VLMs..
We partition the graph $G=(V,E)$ with $|V|$ nodes into three tiers based on the topological centrality of its nodes. Each tier uses a different storage granularity, significantly reducing storage needs while preserving the structural details essential for downstream tasks.
Tier 1: Core Nodes.
This tier consists of nodes with the highest structural influence, functioning as hubs that govern information flow and shape global graph properties. We employ PageRank, a widely adopted centrality measure that reflects both connectivity and reachability, to identify the top $K_{1}\%$ of nodes as the core nodes. For each core node $v$, its complete 2-hop neighborhood $\mathcal{N}_{2}(v)$ is preserved to capture fine-grained local structural information.
Tier 2: Backbone Nodes.
This tier contains nodes located along intercommunity paths, which are essential for preserving global connectivity. These nodes are selected using betweenness centrality, which quantifies how frequently a node lies on shortest paths between other nodes. The next $K_{2}\%$ of nodes are retained, and for each backbone node $v$, only its 1-hop neighborhood $\mathcal{N}_{1}(v)$ is stored to balance representational fidelity with storage efficiency.
Tier 3: Peripheral Nodes.
The remaining nodes are classified as peripheral, whose structural role is primarily defined through their links to higher-tier nodes. To minimize redundancy, we adopt a conditional storage policy: the 1-hop neighborhood $\mathcal{N}_{1}(v)$ of a peripheral node is stored only if it is not connected to any Tier 1 or Tier 2 node. Otherwise, its structural information is implicitly covered by the neighborhoods of higher-tier nodes, ensuring maximal storage efficiency [^3] In practice, $\mathcal{K}$ can be configured to utilize only a subset of tiers (e.g., only Tier 1) to adapt to storage or computational constraints, providing flexibility across application scenarios..
$\mathcal{K}$ not only addresses scalability but also establishes the information backbone for subsequent multimodal collaborative reasoning. In the text modality-based branch, it enables efficient retrieval of relevant structured context, while in the visual modality-based branch, it guides the precise extraction of local visual subgraphs.
3.3 Planning Agent-based Task Routing
This module employs a planning agent to allocate each task to the optimal reasoning modality, either textual or visual. The routing procedure comprises two stages. In the Task Parsing stage, the agent analyzes the natural language question to extract the task intent and key entities. In the Task Categorization stage, the parsed task is classified and dispatched to the corresponding reasoning module.
Task Parsing
Given a natural language question $Q$ over $G$, the planning agent first performs semantic parsing to identify the task type $T$ and extract the set of key entities $\mathcal{E}$:
$(T,\mathcal{E})=f_{\text{parse}}(Q,G)$ (1)
, where $f_{\text{parse}}$ denotes a semantic parsing function implemented by the planning agent. The agent applies few-shot prompting to infer task semantics, producing a task type $T$ (e.g., neighbor identification) and a set of key entity pairs $\mathcal{E}={(v_{i},v_{j})\mid v_{i},v_{j}\in V}$ referenced in the question. The task type $T$ determines the routing strategy, while the entity set $\mathcal{E}$ provides essential context for downstream modules. To ensure uniform processing, queries involving a single entity (e.g., “What is the degree of node $v_{i}$?”) are reformulated as $(v_{i},\emptyset)$ by introducing a null entity placeholder.
Task Categorization and Execution Planning
Following parsing, the planning agent routes the task to the appropriate module based on $T$ and formulates a high-level execution plan [^4] The planning agent first matches the question against predefined task templates. If no match is found, it falls back to semantic parsing to infer the intent and choose the reasoning modality.. Building on the task definitions established in existing graph understanding benchmarks Chen et al. (2024a); Yuan et al. (2025a); Cao et al. (2025), we propose a task classification framework that leverages the complementary strengths of different graph representation modalities. Specifically, graph tasks are categorized into two classes, each handled by a dedicated module:
Simple Property Task Examples: “What is the total number of nodes in this Graph?”; “Return the degree of node A.” Complex Local Task Examples: “Find the shortest path between nodes A and B.”; “Do nodes A and B share any third-order neighbors?”
Simple Property Tasks.
These tasks involve questioning either macroscopic graph statistics or microscopic, explicitly defined properties. Such information is efficiently and reliably obtained from textual graph descriptions, whereas visual representations often suffer from resolution limitations that obscure fine-grained details. Accordingly, the planning agent routes these tasks to the GraphRAG agent module, providing a high-level retrieval strategy for execution over $\mathcal{K}$.
Complex Local Tasks.
These tasks require reasoning over fine-grained topological structures, such as finding the shortest path between nodes. For such tasks, a visual subgraph offers a concise and intuitive representation that is more effective for complex topological reasoning than verbose textual descriptions. Consequently, the planning agent delegates such tasks to the Visual Graph Thoughts agent, while decomposing the task into algorithm-guided substeps and constructing a high-level reasoning plan.
3.4 Visual Modality-based Graph Understanding
This module leverages visual-modality graph representations to handle complex local problems that require fine-grained topological understanding and multi-step logical inference.
To address the scalability issue of large graphs, we extract a task-relevant $k$-hop subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with a maximum of $N_{max}$ nodes from $G$ in $\mathcal{K}$, centered on the key entities $\mathcal{E}$ (details are provided in Appendix B). The resulting subgraph is then transformed into a high-resolution visual representation $G_{\text{image}}=f_{\mathrm{viz}}(G^{\prime})$, which preserves the essential topological structure of $G^{\prime}$ and serves as a faithful basis for downstream reasoning.
3.4.1 Visual Graph Thoughts
Existing text-based CoT methods often suffer from pseudo-visual reasoning, in which visual information gradually fades as the reasoning chain lengthens Chen et al. (2024b); Zhang et al. (2024b); Zou et al. (2024); Cheng et al. (2025). To overcome this limitation in multimodal reasoning for complex local tasks, we introduce Visual Graph Thoughts.
After obtaining the visual subgraph $G_{\text{image}}$, the visual reasoning agent $\mathcal{M}_{\text{VRA}}$, instantiated from a VLM, performs iterative multimodal reasoning under the guidance of a high-level plan $\Pi=(\pi_{1},\pi_{2},\dots,\pi_{n})$ generated by the planning agent.
We formalize this process as a state-transition system. At step $t$ ($1\leq t\leq n$), the reasoning state $S_{t}$ is defined by the current visual representation $G_{\text{image}}^{(t-1)}$, the accumulated reasoning history $H_{t-1}$, and the current plan instruction $\pi_{t}$. Conditioned on $S_{t}$, $\mathcal{M}_{\text{VRA}}$ produces an intermediate reasoning output $o_{t}$ and an action $a_{t}$:
$(o_{t},a_{t})=\mathcal{M}_{\text{VRA}}(S_{t})=\mathcal{M}_{\text{VRA}}(G_{\text{image}}^{(t-1)},H_{t-1},{\pi}_{t})$ (2)
If $a_{t}$ invokes $f_{\mathrm{viz}}$ (e.g., node highlighting), the visual representation is updated as $G_{\text{image}}^{(t)}=f_{\mathrm{viz}}(G_{\text{image}}^{(t-1)},a_{t})$; otherwise, $G_{\text{image}}^{(t)}=G_{\text{image}}^{(t-1)}$. This closed-loop process continues until all plan steps are executed.
Finally, $\mathcal{M}_{\text{VRA}}$ aggregates the reasoning trace $H_{n}$ to produce the final answer $A_{\text{final}}=f_{\text{summarize}}(H_{n})$, where $f_{\text{summarize}}$ aggregates intermediate reasoning outputs to generate the final prediction.
3.4.2 Aligning Visual Graph Thoughts with Process-level DPO
To further improve the multimodal reasoning capability of $\mathcal{M}_{\text{VRA}}$ by promoting the generation of reliable visual graph thoughts, we adopt a process-level Direct Preference Optimization (DPO) training strategy Rafailov et al. (2023). Compared to outcome-based methods Kong et al. (2025a); Liu et al. (2025), we model the entire multimodal reasoning trajectory, which provides stable, step-level reinforcement for graph inference tasks.
1. Construction of Process Preference Dataset.
We construct a process preference dataset $\mathcal{D}$, where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$. Here, $y_{w}$ denotes the “Chosen” path representing the preferred reasoning trajectory, while $y_{l}$ denotes the “Rejected” path representing a non-preferred trajectory. The construction details are in Appendix A.
“Chosen” Path ($y_{w}$).
We construct standardized reasoning templates aligned with the logical structure of each graph task. Based on these templates, we generate ground-truth reasoning steps and answers using large-scale synthetic graphs (e.g., ER and BA graphs Li et al. (2005)) and validate their correctness through graph analysis libraries.
“Rejected” Path ($y_{l}$).
To help the VLM distinguish between preferred and non-preferred paths, we construct negative samples $y_{l}$ by modifying $y_{w}$. Specifically, we extend error construction strategies from existing benchmarks Chen et al. (2024a); Yuan et al. (2025a) to multimodal reasoning tasks. The error categories are summarized in Table 5.
2. DPO Training Process.
To align the generation of visual graph thoughts with ground-truth multimodal reasoning trajectories, we fine-tune $\mathcal{M}_{\text{VRA}}$ using DPO on $\mathcal{D}$. For stability, we use a reference policy $\pi_{\text{ref}}$, defined as a copy of $\mathcal{M}_{\text{VRA}}$ that has been supervised to fine-tune in $\mathcal{D}$ and kept fixed during DPO training. The DPO objective minimizes the following loss function:
$\mathcal{L}_{\mathrm{DPO}}=-\mathbb{E}_{(\mathbf{x},y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\!\left(\beta\log\frac{r_{\theta}(y_{w}|\mathbf{x})}{r_{\theta}(y_{l}|\mathbf{x})}\right)\right]$ (3)
, where the policy ratio is defined as $r_{\theta}(y|\mathbf{x})=\pi_{\theta}(y|\mathbf{x})/\pi_{\mathrm{ref}}(y|\mathbf{x})$, and $\beta$ controls the regularization strength. By maximizing the log-probability margin between the preferred ($y_{w}$) and rejected ($y_{l}$) trajectories, the model is guided away from predefined reasoning errors. The resulting policy generates visual graph thoughts that are both logically coherent and tightly grounded in visual graph evidence throughout the inference process.
3.5 Text Modality-based Graph Understanding
For simple property tasks, we adopt GraphRAG-based retrieval for targeted information extraction. Note that existing RAG variants and optimizations Deng et al. (2023); Lee et al. (2025); Gao et al. (2025); Chang et al. (2025) are orthogonal to GraphVista and can be easily integrated to improve retrieval quality. To maintain generality, we employ the standard RAG formulation in this work.
For a given task request $(Q,\mathcal{E})$, we first employs a retrieval function $f_{\text{retrieve}}$ to question $\mathcal{K}$ and extract relevant context, formalized as $C_{\text{retrieved}}=f_{\text{retrieve}}(\mathcal{K},(Q,\mathcal{E}))$. Subsequently, the downstream VLM is prompted with both the original question $Q$ and the retrieved context $C_{\text{retrieved}}$ to generate the final answer.
Table 1: Performance comparison on Grena and GraphSQA Benchmark. Red and Blue highlight the best and second-best results respectively.
Method Model Grena GraphSQA
Simple Complex Overall Simple Complex Overall
Text-only InternVL3-9B 0.1211 0.0287 0.0829 0.4788 0.2498 0.3169
GLM-4.1V-9B 0.0157 0.0071 0.0121 0.2532 0.1670 0.1929
Qwen2.5-VL-7B 0.0096 0.0095 0.0096 0.2340 0.1122 0.1488
Qwen3-8B 0.0065 0.0056 0.0061 0.1351 0.1352 0.1352
Gemma-3-12B 0.0369 0.0358 0.0364 0.5917 0.3245 0.4033
GraphPRM (DPO) Qwen2.5-VL-7B (DPO) 0.0332 0.0087 0.0231 0.7198 0.3455 0.4550
Qwen3-8B (DPO) 0.0060 0.0049 0.0055 0.3629 0.2357 0.2728
Gemma-3-12B (DPO) 0.0417 0.0335 0.0383 0.6235 0.3455 0.4195
GraphInsight InternVL3-9B 0.1788 0.0056 0.1073 0.7564 0.2315 0.3860
GLM-4.1V-9B 0.1849 0.0016 0.1092 0.4313 0.3273 0.3583
Qwen2.5-VL-7B 0.1925 0.0085 0.1165 0.5769 0.2293 0.3348
Gemma-3-12B 0.2919 0.0236 0.1811 0.7372 0.2478 0.3779
GraphToken Qwen2.5-VL-7B 0.0981 0.0782 0.0899 0.0388 0.1770 0.1376
Qwen3-8B 0.1589 0.0757 0.1245 0.1742 0.1508 0.1579
Gemma-3-12B 0.0158 0.0576 0.0331 0.0566 0.1091 0.0927
GITA InternVL3-9B 0.1751 0.1098 0.1481 0.3910 0.1755 0.2402
GLM-4.1V-9B 0.1651 0.1175 0.1454 0.3397 0.1562 0.2112
Qwen2.5-VL-7B 0.1351 0.0632 0.1054 0.3205 0.1596 0.2079
Gemma-3-12B 0.1732 0.1435 0.1609 0.2692 0.1739 0.2026
GraphVista (Ours) InternVL3-9B 0.9357 0.3578 0.6970 0.7970 0.6886 0.7211
GLM-4.1V-9B 0.9387 0.2727 0.6637 0.7954 0.6657 0.7046
Gemma-3-12B 0.9362 0.3477 0.6932 0.7991 0.6815 0.7168
Gemma-3-4B (DPO) 0.9362 0.3397 0.6899 0.7943 0.6780 0.7129
Qwen2.5-VL-7B (DPO) 0.9367 0.3912 0.7114 0.7973 0.6806 0.7156
4 Evaluation
4.1 Experimental Setup
4.1.1 Baselines
We compare GraphVista with representative state-of-the-art baselines across three categories: (a) Text-based methods, which reason solely over textual graph descriptions, including the standard Text-only approach, GraphPRM Peng et al. (2025) trained via DPO, and GraphInsight Cao et al. (2025). (b) GNN-based methods, which encode structural information into latent representations using graph neural networks. A representative example is GraphToken [^5] These methods are orthogonal to GraphVista, focusing on graph encoding rather than improving the intrinsic graph understanding of VLMs. We include them for completeness. Perozzi et al. (2024). (c) Hybrid methods, exemplified by GITA Wei et al. (2024), which integrate both visual and textual graph modalities into VLMs.
4.1.2 Benchmarks and Evaluation Tasks
Existing graph understanding benchmarks suffer from limited coverage of graph scales and topologies, a lack of task categorization grounded in fundamental graph properties, and insufficient support for evaluating multi-step visual reasoning. To address these gaps, we introduce Grena, a large-scale, multi-task benchmark we designed to evaluate graph understanding of LLMs and VLMs, which encompasses a wide range of task types found in existing benchmarks Wang et al. (2023); Yuan et al. (2025a); Zhu et al. (2025).
Grena spans a wide diversity of graph scales, with node counts ranging up to 2050, and encompasses 18 distinct task types, which are organized into two categories: (1) 7 Simple Property Tasks (Simple) focus on querying basic properties and local structures, and (2) 11 Complex Local Tasks (Complex) require fine-grained topological information and performing fine-grained inference. Further details are provided in Appendix A.
In our experiments, we use Grena to evaluate graph understanding performance on large-scale graphs ($|V|\in[50,2050]$), covering a total of 13,860 tasks, which are split into 8,136 simple and 5,724 complex instances. Complementarily, we use the GraphSQA benchmark Cao et al. (2025) for small-graph understanding, with $|V|$ in the range of $[15,50]$. Across all tasks and benchmarks, we use Accuracy (ACC) as the evaluation metric. Methods that require DPO are trained on $\mathcal{D}$.
(a) Complex (Gemma-3)
(b) Simple (Gemma-3)
(c) Complex (Qwen2.5-VL)
(d) Simple (Qwen2.5-VL)
Figure 3: Analysis on Graphs with Different $|V|$
Table 2: Ablation study on reasoning strategies of GraphVista. The best results within each VLM group are bolded, and the second-best results are underlined.
Strategy Simple Complex Overall
InternVL3-9B
CoT (Baseline) 0.9368 0.1422 0.6086
Visual Graph Thoughts (Text) 0.2220 0.6416
Visual Graph Thoughts (Visual) 0.3578 0.6977
Gemma-3-12B
CoT (Baseline) 0.9362 0.1406 0.6076
Visual Graph Thoughts (Text) 0.2447 0.6506
Visual Graph Thoughts (Visual) 0.3477 0.6932
GLM-4.1V-9B
CoT (Baseline) 0.9387 0.1406 0.6091
Visual Graph Thoughts (Text) 0.2300 0.6460
Visual Graph Thoughts (Visual) 0.2727 0.6637
Qwen2.5-VL-7B (DPO)
CoT (Baseline) 0.9367 0.1414 0.6083
Visual Graph Thoughts (Text) 0.3429 0.6915
Visual Graph Thoughts (Visual) 0.3912 0.7114
4.1.3 Models
We instantiate GraphVista and all baselines using recent open-source, competitive VLMs, including Gemma-3-12B, Gemma-3-4B Kamath et al. (2025), Qwen2.5-VL-7B Bai et al. (2025), InternVL3-9B Zhu and et al (2025), and GLM-4.1V-9B-Thinking Hong et al. (2025), a reasoning-oriented VLM. For text-based methods, we include more capable LLMs such as Qwen3-8BYang et al. (2025) to ensure a fairer comparison.
4.2 Performance
As shown in Table 1, GraphVista significantly outperforms all baseline methods on both Grena and GraphSQA benchmarks.
4.2.1 Simple Property Tasks
[…]
4.2.2 Complex Local Tasks
[…]
Table 3: Ablation results comparing frozen base models (“Frozen”) with models further optimized via DPO (“DPO”).Red and Blue highlight the best and second-best results respectively.
Model Strategy Simple Complex Overall
Gemma-3-4B Frozen 0.9362 0.2933 0.6707
DPO 0.3397 0.6899
Qwen2.5-VL-7B Frozen 0.9367 0.3240 0.6837
DPO 0.3912 0.7114
4.3 Ablation Study
4.3.1 Effectiveness of Reasoning Strategies
We evaluate the effect of visual grounding by comparing Visual Graph Thoughts (Visual) with its text-only variant and the standard textual CoT, where visual subgraphs are replaced by textual descriptions. In Table 2, the visual variant consistently outperforms the baselines, especially on complex tasks, achieving up to $1.1\times$ and $2.8\times$ higher accuracy than the text-only and CoT settings, respectively. These highlight the importance of explicit structural grounding, as purely textual reasoning often yields ambiguous interpretations.
4.3.2 Impact of DPO
We evaluate the impact of process-level DPO by comparing fine-tuned models with their frozen-base counterparts. As reported in Table 3, DPO enhances performance across all settings, with the most significant gains on complex tasks. For instance, DPO boosts the accuracy of Qwen2.5-VL-7B on complex tasks from 0.3240 to 0.3912. These results validate that aligning model generation with expert visual reasoning trajectories is highly effective for complex local tasks.
(a) Gemma-3 (b) Qwen2.5-VL
Figure 4: Hyperparameter analysis on $k$ and $N_{max}$
Figure 5: Hyperparameter analysis on storage strategy of $\mathcal{K}$
4.4 Hyperparameter Analysis
4.4.1 Subgraph Extraction Parameters $k$ and $N_{max}$
We analyze the impact of $k$ and $N_{max}$ on performance for complex local tasks. As shown in Figure 4, Gemma-3 reaches its highest accuracy with a 2-hop neighborhood, indicating that it can leverage a broader structural context. In contrast, Qwen2.5-VL performs best with a 1-hop neighborhood, suggesting that larger subgraphs may introduce extraneous information. These observations point to the need for adapting subgraph extraction strategies to the characteristics of different VLM architectures.
4.4.2 Storage Strategy of $\mathcal{K}$
We analyze the effect of pruning node tiers in $\mathcal{K}$, which are defined by the parameters $K_{1}$% and $K_{2}$%. As demonstrated in Figure 5, performance on simple property tasks improves as more tiers of the graph are stored. Retaining just the top $K_{1}$% of nodes (Tier 1) along with their 2-hop neighborhoods already provides a strong baseline for localized retrieval. Adding the next $K_{2}$% of nodes (Tier 2) enhances performance further, although the gains diminish when including peripheral nodes (Tier 3). While full graph storage is essential for precise macroscopic tasks where any pruning leads to information loss, these findings reveal a clear trade-off between performance and storage efficiency for general reasoning tasks.
5 Conclusion
In this paper, we propose GraphVista, a unified framework addressing scalability and modality coordination in graph understanding. To improve scalability, GraphVista uses a lightweight hierarchical GraphRAG base to compress and store task-relevant information. For modality coordination, the planning agent routes tasks to the most suitable modality, text modality for simple property tasks or visual modality for complex local tasks. Future work will extend GraphVista to more complex graph types with labels and semantics.
Limitations
The proposed GraphVista framework offers a promising approach for scalable, modality-coordinated graph understanding, while also presenting several potential extensions that could inspire future research directions.
First, while the framework currently focuses on graph topology, its modular design is intended to facilitate extensions for handling rich semantic information. For instance, a semantic extraction module could be easily incorporated to process node attributes and edge labels. This would broaden the framework’s applicability to domains such as knowledge graph reasoning and question answering over heterogeneous networks. Additionally, for the lightweight GraphRAG base, we adopted a general and efficient implementation to ensure the framework’s versatility. We note that many advanced optimization strategies, such as sophisticated indexing and retrieval techniques, are orthogonal to our core contributions. A key advantage of GraphVista’s modular design is that these specialized modules can be readily integrated into our framework. Therefore, future work could easily incorporate state-of-the-art components to further enhance the performance of our framework, demonstrating its flexibility.
Appendix A Grena Benchmark
Existing graph understanding benchmarks suffer from three major limitations: (i) limited coverage of graph scales and topological structures; (ii) lack of task categorization grounded in fundamental graph properties; and (iii) insufficient support for modeling and evaluating multi-step visual reasoning in graph inference.
To address these gaps, we propose the Grena Benchmark, a large-scale, multi-task benchmark with step-level visual reasoning capability. It is designed to rigorously evaluate and advance the cognitive and reasoning abilities of Vision-language models (VLMs) in graph domains.
The core design of Grena Benchmark is reflected in the following three aspects:
Table 5: Categories of errors used to construct $y_{l}$.
Textual and Logical Errors
Factual Errors Incorrect identification of a neighbor set $\mathcal{N}(v)$ or an edge weight $w(e)$.
Logical Errors Faults in algorithmic reasoning (e.g., using edge count instead of weight in a shortest-path task).
Computation Errors Mistakes in numerical operations such as counting.
Omitted Steps Skipping essential reasoning steps.
Multimodal Errors
Element Misrecognition Misidentifying nodes or edges in the visual representation.
Visual Neglect Failure to incorporate visual information, with reasoning based only on text.
Text–Visual Inconsistency Mismatch between textual reasoning and visual grounding.
Visualization Misuse Incorrect or missing calls to visualization functions.
A.1 Scalability and Structural Diversity
A primary objective of Grena is to systematically evaluate the scalability graph understanding capabilities of VLMs. To this end, the benchmark includes graphs that span a wide range of scales, with node counts from 50 to 2,050. This diversity in size is essential for assessing how model performance varies with increasing graph size and structural complexity.
In addition to scale, Grena incorporates diverse topological structures by including both Erdős–Rényi (ER) random graphs and Barabási–Albert (BA) scale-free networks. While ER graphs provide a baseline for performance on uniform structures, BA graphs feature power-law degree distributions and hubs, closely resembling the topology of many real-world systems. This combination of scales and topologies creates a robust testbed for evaluating model generalization across heterogeneous graph structures.
A.2 Categorization of Graph Understanding Tasks
To provide a structured assessment of reasoning abilities, Grena Benchmark organizes tasks into two categories:
  • Simple QA Tasks. These tasks focus on querying basic properties and local structures of graphs, such as “What is the total number of nodes?” or “What is the degree of node $v_{A}$?” They primarily evaluate a model’s ability to store and retrieve graph information. With a highly regularized form, these tasks emphasize breadth of coverage and scalability.
  • Complex Reasoning Tasks. These tasks require models to integrate fine-grained topological structures and perform logical reasoning. For example, shortest-path problems demand identifying the optimal path from node $v_{A}$ to node $v_{B}$. Compared with retrieval tasks, these tasks probe deeper reasoning capabilities by requiring structural integration and inference.
A.3 Process-Level Supervision for Visual Reasoning
A key contribution of the Grena Benchmark is its native support for training models on step-level visual reasoning. For tasks that require a chain-of-thought process (e.g., pathfinding), the benchmark provides not only the final answers but also intermediate visual exemplars. These include the original graph and a series of "state graphs" that progressively highlight key nodes, edges, or subgraphs. This design provides dense supervisory signals, making it suitable for both Supervised Fine-Tuning (SFT) and alignment techniques like Direct Preference Optimization (DPO), thereby enabling VLMs to learn explicit reasoning processes.
To facilitate these training paradigms, we construct a high-quality, process-centric dataset where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$, representing the input, a preferred reasoning path (Chosen), and a non-preferred one (Rejected):
  • Supervised Fine-Tuning Data (The "Chosen" Path, $y_{w}$). The "Chosen" paths serve as gold-standard examples for SFT. They are generated based on human-annotated, standardized reasoning templates that align with the logical flow of each task. We use large-scale synthetic graphs (e.g., Erdős–Rényi and Barabási–Albert models), and programmatically validate the correctness and completeness of all reasoning steps using libraries like NetworkX.
  • Direct Preference Optimization Data (The "Rejected" Path, $y_{l}$). To create preference pairs $(y_{w},y_{l})$ for DPO, the "Rejected" paths are systematically generated by introducing errors into the "Chosen" paths. These error-construction strategies are adapted from established textual reasoning benchmarks and extended to multimodal contexts. The resulting paths feature common reasoning pitfalls (as detailed in Table 5), teaching the model to distinguish between correct and flawed reasoning.
Table 6: Task types in the Grena Benchmark: definition, expected output and category.
Task Name Definition and Expected Output
Bipartite Detection Determine whether $G=(V,E)$ is bipartite (2-colorable). Output: boolean (is bipartite?), optionally a bipartition $(S,T)$.
Clique Detection Detect the existence of (or find) a clique of specified size $k$, or return maximal cliques. Output: boolean (exists $k$-clique?) or node set(s) forming a clique.
Common Third-Order Neighbors Given nodes $u,v$, compute the intersection $N^{(3)}(u)\cap N^{(3)}(v)$ where $N^{(3)}(\cdot)$ denotes nodes at distance exactly 3. Output: set of nodes.
Connectivity Check Decide whether $G$ is connected (single connected component). Output: boolean.
Connectivity Analysis Full analysis of connectivity: number of connected components, sizes of components, and component ID/size for a query node. Output: integer(s) and list(s) (component counts and sizes).
Critical Node Detection Identify articulation points (nodes whose removal increases the number of connected components) or nodes critical to connectivity. Output: set of node ids.
Cycle Detection Determine whether any cycle exists; optionally return an example cycle (sequence of nodes). Output: boolean and, if requested, a cycle path.
Edge Count Return the total number of edges $|E|$. Output: integer.
Edge Existence Given a pair $(u,v)$, decide whether edge $(u,v)\in E$. Output: boolean.
Find Connected Edges Given a node $v$ (or a small subgraph), return the set of incident edges or the edges that connect the queried substructure. Output: set of edges (pairs).
Highest-Degree Neighbor For a query node $v$, return the neighbor $w\in N(v)$ with the largest degree $\deg(w)$ (tie-breaking rule: smallest id or deterministic rule). Output: node id.
Neighbor Connections Measure connections among the neighbors of $v$: e.g., number of edges in the induced subgraph $G[N(v)]$. Output: integer (edge count) or list of edges.
Node Count Return the total number of nodes $|V|$. Output: integer.
Node Degree Given node $v$, return its degree $\deg(v)$ (or weighted degree if weights provided). Output: integer (or numeric).
Planarity Testing Determine whether $G$ is planar; optionally, produce a planar embedding if planar. Output: boolean and (optional) embedding.
Shortest Path Given nodes $u,v$, compute a shortest path (and its length). Supports unweighted or weighted variants. Output: sequence of nodes (path) and length.
Third-Order Neighbors For node $v$, return $N^{(3)}(v)$, the set of nodes at graph distance exactly 3 from $v$. Output: set of nodes.
Triangle Counting Count all triangles (3-cycles) in $G$, and/or enumerate them. Output: integer (triangle count) or list of 3-tuples (node triples).
Appendix B Subgraph Extraction
To mitigate the scalability challenges of large graphs, we first extract a task-relevant local subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ from $G$ stored in $\mathcal{K}$, based on the key entities $\mathcal{E}$, formalized as:
$G^{\prime}=f_{\text{extract}}(G,\mathcal{E})$ (4)
, where the extraction strategy $f_{\text{extract}}$ is determined by the structural characteristics of $\mathcal{E}$, and can be categorized as follows:
Ego-centric
For problems involving a single entity, where $\mathcal{E}=(v_{i},\emptyset)$, we extract its $k$-hop neighborhood subgraph $G^{\prime}_{v_{i},k}=(V^{\prime},E^{\prime})$. The node set $V^{\prime}$ is defined as:
$V^{\prime}=\{v\in V\mid\text{dist}_{G}(v,v_{i})\leq k\}$ (5)
, where $\text{dist}_{G}(u,v)$ denotes the shortest path distance between nodes $u$ and $v$ in $G$. The edge set $E^{\prime}$ is induced from $G$ by $V^{\prime}$.
Multi-centric
For problems concerning relationships between multiple entities, where $\mathcal{E}=(v_{s},v_{t})$, we extract paths connecting the source $v_{s}$ and target $v_{t}$. We employ Yen’s K-shortest paths to obtain $K$ candidate paths $\mathcal{P}=\{P_{1},\dots,P_{K}\}$. We then construct a set $V_{\mathcal{P}}=\bigcup_{j=1}^{K}V(P_{j})$ of all nodes on these paths. Finally, a 1-hop neighborhood expansion is performed on all nodes in $V_{\mathcal{P}}$ to form the final subgraph node set $V^{\prime}=\bigcup_{v\in V_{\mathcal{P}}}\{u\in V\mid\text{dist}_{G}(u,v)\leq 1\}$. The edge set $E^{\prime}$ is induced by $V^{\prime}$ from $G$.
Appendix C Subgraph Pruning Algorithms
To maintain visual clarity and adhere to computational constraints, particularly for large graphs, the size of an extracted subgraph $G^{\prime}$ is bounded by a maximum node count, $N_{max}$. In instances where the initial extracted subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ exceeds this threshold, a principled pruning procedure is invoked to reduce its size while preserving the most topologically significant nodes. The specific pruning strategy is contingent upon the extraction method used, namely ego-centric or multi-centric, as detailed in Algorithm 1 and Algorithm 2, respectively.
C.1 Ego-centric Subgraph Pruning
This approach, detailed in Algorithm 1, addresses subgraphs generated around a single central node, $v_{i}$. The core principle is to retain nodes most relevant to this central entity. All nodes except for $v_{i}$ are considered candidates for removal and are sorted into a priority list for pruning based on the hierarchical key $(-d(v,v_{i}),t(v),c(v))$. This multi-level sorting criterion prioritizes the removal of nodes based on the following order:
  1. 1.
    Distance: Nodes farthest from the central node $v_{i}$ (descending distance $d$) are removed first.
  2. 2.
    Structural Tier: Among nodes at the same distance, those in lower-importance tiers within the Hierarchical GraphRAG Base $\mathcal{K}$ (ascending tier $t(v)$) are prioritized for removal.
  3. 3.
    Centrality: For nodes with the same distance and tier, those with a lower centrality score (ascending centrality $c(v)$) are removed.
By systematically applying these criteria, the algorithm ensures that the pruned subgraph retains the most structurally-important neighborhood around the central node.
C.2 Multi-centric Subgraph Pruning
This strategy, presented in Algorithm 2, is designed for subgraphs constructed around a set of K-shortest paths, $V_{\mathcal{P}}$, connecting multiple entities. The primary objective is to preserve the integrity of these critical paths. Therefore, all nodes lying on these paths ($V_{\mathcal{P}}$) are explicitly protected from pruning. Candidate nodes for removal are those not on the primary paths, and their sorting priority is determined by the key $(t(v),c(v))$, which prioritizes removal based on:
  1. 1.
    Structural Tier: Nodes in lower-importance tiers (ascending tier $t(v)$).
  2. 2.
    Centrality: Among nodes in the same tier, those with a lower centrality score (ascending centrality $c(v)$).
This approach effectively reduces the subgraph’s complexity while guaranteeing that the core relational structure between the key entities remains intact.
Input : $G^{\prime}=(V^{\prime},E^{\prime})$: Initial $k$-hop neighborhood subgraph.
$v_{i}$: Central node for extraction.
$N_{\max}$: Maximum allowed number of nodes.
$\mathcal{K}$: Hierarchical GraphRAG Base.
Output : $G^{\prime}_{\text{pruned}}$: Pruned subgraph.
1ex
if $|V^{\prime}|\leq N_{\max}$ then return $G^{\prime}$
$V_{\text{cand}}\leftarrow V^{\prime}\setminus\{v_{i}\}$
// Sort by priority: distance $\downarrow$, tier $\uparrow$, centrality $\uparrow$
Sort $V_{\text{cand}}$ into list $L$ by key $(-d(v,v_{i}),t(v),c(v))$
$n_{\text{prune}}\leftarrow|V^{\prime}|-N_{\max}$
$V_{\text{prune}}\leftarrow L[1:n_{\text{prune}}]$
$V_{\text{kept}}\leftarrow V^{\prime}\setminus V_{\text{prune}}$
$G^{\prime}_{\text{pruned}}\leftarrow$ subgraph of $G^{\prime}$ induced by $V_{\text{kept}}$
return $G^{\prime}_{\text{pruned}}$
Algorithm 1 Ego-centric Subgraph Pruning
Input : $G^{\prime}=(V^{\prime},E^{\prime})$: Subgraph expanded from $K$-shortest paths.
$V_{\mathcal{P}}$: Nodes on the $K$-shortest paths.
$N_{\max}$: Maximum allowed number of nodes.
$\mathcal{K}$: Hierarchical GraphRAG Base.
Output : $G^{\prime}_{\text{pruned}}$: Pruned subgraph.
1ex
if $|V^{\prime}|\leq N_{\max}$ then return $G^{\prime}$
$V_{\text{cand}}\leftarrow V^{\prime}\setminus V_{\mathcal{P}}$
// Sort by priority: tier $\uparrow$, centrality $\uparrow$
Sort $V_{\text{cand}}$ into list $L$ by key $(t(v),c(v))$
$n_{\text{prune}}\leftarrow|V^{\prime}|-N_{\max}$
$V_{\text{prune}}\leftarrow L[1:n_{\text{prune}}]$
$V_{\text{kept}}\leftarrow V^{\prime}\setminus V_{\text{prune}}$
$G^{\prime}_{\text{pruned}}\leftarrow$ subgraph of $G^{\prime}$ induced by $V_{\text{kept}}$
return $G^{\prime}_{\text{pruned}}$
Algorithm 2 Multi-centric Subgraph Pruning
Table 7: Key Hyperparameters and Framework Configurations.
Category Parameter Value
Model & Inference Configuration
Model Path MODEL_PATH (Path to VLM)
Tensor Parallel TENSOR_PARALLEL_SIZE 1
GPU Memory Util. GPU_MEMORY_UTILIZATION 0.4
Max Model Length MAX_MODEL_LEN 4096
Generation Sampling Parameters
Temperature TEMPERATURE 0.01
Max Tokens MAX_TOKENS 2048
Top P TOP_P 0.9
Subgraph Extraction
Max Subgraph Nodes max_nodes 25
Max Hops max_hops 2
Appendix D Notations
This section summarizes all notations used throughout this paper, as in Table 8.
Table 8: Summary of notations used in this paper.
Notation Description
$G=(V,E)$ Graph with node set $V$ and edge set $E$.
$\mathcal{K}$ The Hierarchical GraphRAG Base.
$Q,T,\mathcal{E}$ NL question, its task type, and key entities.
$f_{\text{parse}}$ Semantic parsing function for questions.
$\mathcal{N}_{k}(v)$ $k$-hop neighborhood of node $v$.
$|V|,N_{\text{max}}$ Total nodes in graph; max nodes in a subgraph.
$K_{1}\%,K_{2}\%$ Node percentage in Tier 1 (Core) and Tier 2 (Backbone).
$G^{\prime}=(V^{\prime},E^{\prime})$ A subgraph extracted from the original graph $G$.
$f_{\text{viz}},G_{\text{image}}$ Subgraph visualization function and its visual output.
$\mathcal{M}_{\text{VRA}}$ The Visual Reasoning Agent.
$\Pi,S_{t},H_{t}$ Reasoning plan, state at step $t$, and history up to $t$.
$(o_{t},a_{t})$ Intermediate output and action at step $t$.
$C_{\text{retrieved}}$ Retrieved context from the GraphRAG Base $\mathcal{K}$.
$f_{\text{retrieve}}$ Retrieval function for querying $\mathcal{K}$.
$\mathcal{D},(x,y_{w},y_{l})$ Preference dataset and sample (input, chosen, rejected).
$\pi_{\theta},\pi_{\text{ref}}$ DPO policy and fixed reference models.
$r_{\theta}(y|x)$ Ratio of policy probabilities $\pi_{\theta}(y|x)/\pi_{\text{ref}}(y|x)$.
$\beta,\mathcal{L}_{\text{DPO}}$ DPO regularization strength and its loss function.
$f_{\text{extract}}$ The subgraph extraction function.
$dist_{G}(u,v)$ Shortest path distance between nodes $u$ and $v$.
$V_{\mathcal{P}}$ Set of nodes on the K-shortest paths.
$d(v,v_{i}),t(v),c(v)$ Node’s distance to center $v_{i}$, structural tier, and centrality.
Appendix E Experimental Settings
Our experiments are conducted using 8 NVIDIA H20 GPUs. The framework is implemented in Python 3.11 with CUDA Version 12.6, leveraging the vLLM library for efficient inference of the VLMs.
Appendix F Implementation Details
Key hyperparameters and framework configurations are detailed in Table 7.
(a) Original graph
(b) Identify neighbors
(c) Check edges
(d) Triangle confirmed
Figure 6: Visualizing the process of Triangle Detection for node A. The algorithm identifies node A’s neighbors and then checks for edges between them to confirm a 3-cycle.
(a) Identify start (B) and end (F)
(b) Explore from node B
(c) Path extended to node D
(d) Path extended to node E
(e) Shortest path found
Figure 7: Visualizing a breadth-first search for the Shortest Path from node B to F. The algorithm explores neighbors layer-by-layer until reaching the destination.
(a) Step 1: Identify target node A
(b) Step 2: Highlight direct neighbors
Figure 8: Visualizing Neighbor Retrieval for a target node A. The process identifies all nodes that are directly connected to node A.
Appendix G Example of Visual Graph Thought Chains
This section provides concrete examples of the step-by-step "Visual Graph Thoughts" process, a core component of our framework for solving complex local reasoning tasks. Each figure-set deconstructs a graph-based problem into a sequence of visual states. This methodology allows the VLM to perform grounded reasoning by identifying and highlighting relevant nodes and edges at each step, culminating in a final solution.
Figure 6 illustrates the process for Triangle Detection. The chain begins by identifying the neighbors of a central node, then visually inspects for edges between those neighbors to confirm a 3-cycle. For Shortest Path finding, shown in Figure 7, the model emulates a breadth-first search, extending the path layer-by-layer from the source until the destination is reached. Finally, Figure 8 depicts the fundamental operation of Neighbor Retrieval, where the model identifies and highlights all nodes directly adjacent to a given target. These examples demonstrate how Visual Graph Thoughts provide explicit, verifiable reasoning grounded in visual evidence.
Appendix H Prompt and QA Templates
This section provides prompt templates for evaluating the graph understanding capabilities of Vision-Language Models (VLMs) within the GraphVista framework. The templates are categorized into Simple Property Tasks and Complex Local Tasks, reflecting the task routing mechanism of the planning agent as described in Section 3.3 of the main paper. Each template includes an example prompt, the expected model behavior based on the GraphVista architecture, and the rationale behind the chosen modality.
H.1 Simple Property Tasks: Text Modality with GraphRAG
For simple property tasks, the planning agent routes the query to the text-modality branch. This branch utilizes GraphRAG-based retrieval on the Hierarchical GraphRAG Base ($\mathcal{K}$) to efficiently extract factual information without needing visual inference.
H.2 Complex Local Tasks: Visual Modality with Visual Graph Thoughts
For tasks requiring nuanced topological understanding or multi-step reasoning, the planning agent delegates the task to the visual modality branch. This module extracts a relevant subgraph, visualizes it, and employs the Visual Graph Thoughts agent to perform grounded, step-by-step multimodal reasoning.
Task 1: Node Existence and Properties Prompt Template: "Determine if node <node_id> exists in the graph. If so, what is its degree?" GraphVista Behavior: The planning agent categorizes this as a simple property task. The model performs a targeted retrieval query against the GraphRAG Base $\mathcal{K}$ to find the entry for <node_id>. The retrieved context, containing the node’s 1-hop neighborhood information, is then used to generate a direct answer. Rationale: This task involves retrieving explicitly stored attributes. The textual representation in the GraphRAG base is optimized for such direct lookups, making it efficient and reliable.
Task 2: Edge Existence and Weight Prompt Template: "Is there a direct edge between node <node_id_1> and node <node_id_2>? If yes, provide its weight." GraphVista Behavior: The query is identified as a simple property task. The system retrieves the adjacency information for <node_id_1> from $\mathcal{K}$. The VLM then processes this structured text to check for a connection to <node_id_2> and extracts the corresponding weight attribute. Rationale: Like node property lookups, edge verification relies on retrieving specific, pre-processed information. The textual modality avoids the potential for visual ambiguity (e.g., overlapping edges or nodes) that could occur in a complex graph visualization.
Task 3: Shortest Path Identification Prompt Template: "Find and list the sequence of nodes that form the shortest path between node <node_id_1> and node <node_id_2>." GraphVista Behavior: The planning agent categorizes this as a complex local task. The Visual Graph Thoughts Agent extracts a multi-centric subgraph centered around the shortest path candidates between <node_id_1> and <node_id_2>. This high-resolution subgraph is visualized. The VLM then initiates a Visual Graph Thought process, highlighting nodes and edges step-by-step in the image to trace the optimal path, verbalizing its reasoning at each step before presenting the final sequence. Rationale: Textual descriptions of paths can be convoluted and difficult for VLMs to reason over effectively. A visual representation provides a “what-you-see-is-what-you-get” view of the topology, which is more intuitive for complex structural reasoning and naturally supports stepwise inference grounded in visual evidence.
Task 4: Triangle Detection Prompt Template: "Does a 3-cycle (triangle) involving node <node_id> exist? If so, list the nodes of one such triangle." GraphVista Behavior: This is routed to the visual modality. An ego-centric subgraph around <node_id> is extracted from $\mathcal{K}$ and visualized. The Visual Graph Thoughts agent first identifies the neighbors of <node_id> in the image. It then visually inspects for edges connecting any two of these neighbors. If such an edge is found, the VLM confirms the existence of a triangle and outputs the three nodes forming the cycle. Rationale: Detecting local structural patterns like cycles is inherently a topological problem. Visual inspection allows the VLM to directly perceive the connectivity pattern, whereas a text-based approach would require inefficient and error-prone traversal of adjacency lists to check for path closures.

### v2

@@
-See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models
+See or Say Graphs: Agent-Driven Scalable Graph Structure Understanding with Vision-Language Models
 Abstract
 […]
-See or Say Graphs: Agent-Driven Scalable Graph Understanding
-with Vision-Language Models
+See or Say Graphs: Agent-Driven Scalable Graph Structure
+Understanding with Vision-Language Models
 Shuo Han1,2*   Yukun Cao1,2*   Zezhong Ding1,2 Zengyi Gao1,2S.
 Kevin Zhou1,3Xike Xie1,2 1University of Science and Technology of China, China 2Data Darkness Lab, MIRACLE Center, USTC, China 3MIRACLE Center, Suzhou Institute for Advance Research, USTC, China {shuo.han, ykcho, zezhongding, gzy02}@mail.ustc.edu.cn {skevinzhou, xkxie}@ustc.edu.cn
-**footnotetext: These authors contributed equally to this work.footnotetext: Corresponding author.
 1 Introduction
-Recently, vision-language models (VLMs) have shown potential for general-purpose understanding of graph data Tang et al.
+Recently, vision-language models (VLMs) have shown potential for general-purpose graph structure understanding Tang et al.
 (2024a); Kong et al.
-(2025b); Ding et al.
+(2025); Ding et al.
 (2025b), offering new paradigms for solving real-world problems naturally represented as graphs.
@@
 (2025).
-To this end, existing studies have explored two primary graph input modalities for VLMs: textual graph descriptions [^1] Many early studies on text-modality-based graph understanding were developed on LLMs.
-As VLMs extend LLMs with visual capabilities, this paper presents a unified discussion from the perspective of graph understanding with VLMs., which convert structural information into natural language for indirect memory and reasoning over node properties and relations Jiang et al.
+To this end, existing studies have explored two primary graph input modalities for VLMs: textual graph descriptions [^1] Most early studies Cao et al.
+(2025); Guan et al.
+(2025); Peng et al.
+(2025) on text-modality-based graph understanding were developed on LLMs.
+As VLMs extend LLMs with visual capabilities, this paper presents a unified discussion from the perspective of graph structure understanding with VLMs., which convert structural information into natural language for indirect memory and reasoning over node properties and relations Jiang et al.
 (2023); Jin et al.
@@
 (2024c); Zhu et al.
-(2025).
+(2025b).
 More recent efforts further combine these modalities to improve structural understanding and complex reasoning on graphs Wei et al.
 (2024).
-Figure 1: Comparison of textual and visual modalities across different types of graph understanding tasks
-However, despite these advances, current VLMs still lack a unified framework that systematically addresses the scalability challenges and the difficulties in coordinating textual and visual modalities inherent in graph understanding.
+Figure 1: Comparison of textual and visual modalities across graph structure understanding tasks
+However, despite these, current VLMs still lack a unified framework that systematically addresses the scalability challenges and the difficulties in coordinating textual and visual modalities inherent in graph structure understanding.
 Specifically:
@@
 Visual graph inputs, on the other hand, are restricted by image resolution, resulting in the loss of structural details as graph size increases Zhu et al.
-(2025).
-In addition, VLMs face the lack of an effective mechanism to coordinate different input modalities for graph understanding, which prevents them from fully leveraging the distinct strengths and complementary advantages of textual and visual modalities.
-As illustrated in Figure 1, under the token length limits of VLMs, the text modality is more effective for simple property tasks such as counting nodes, measuring degrees, and assessing connectivity.
+(2025b).
+In addition, VLMs face the lack of an effective mechanism to coordinate different input modalities for graph structure understanding, which prevents them from fully leveraging the distinct strengths and complementary advantages of textual and visual modalities.
+As illustrated in Figure 1, under the token length limits of VLMs, the text modality is more effective for tasks that require direct statistical access to explicit graph properties, such as counting nodes, measuring degrees, and assessing connectivity.
 In these cases, the information can be directly located and reasoned over with minimal effort.
 By contrast, the visual modality often performs worse because the limited resolution of global visual graphs blurs structural details, causing VLMs to misidentify relevant elements and leading to performance degradation.
-For tasks that require local and structurally complex reasoning, such as shortest-path computation or cycle detection, high-resolution local visual subgraphs offer a clear advantage for VLMs.
-They provide a “what-you-see-is-what-you-get” view of the topology and naturally support stepwise reasoning.
+For tasks that require local graph structure reasoning [^2] Tasks based on global graph structure require full-graph access coupled with fine-grained structural reasoning, and can be viewed as a composition of global information retrieval and local structural-reasoning tasks., such as shortest-path computation or cycle detection, high-resolution local visual subgraphs offer a clear advantage for VLMs.
+They provide a “what-you-see-is-what-you-get” view of the graph topology and naturally support structural reasoning.
 In contrast, textual descriptions are often lengthy and lack structural clarity, leaving VLMs with only indirect language cues and thus little structural grounding for effective complex graph reasoning.
-Therefore, a natural conclusion is that realizing the full potential of VLMs in graph understanding requires addressing these two challenges.
-This entails, on the one hand, preserving as much task-relevant graph information as possible within limited input tokens, and on the other hand, assigning tasks according to modality strengths: letting the text modality efficiently address simple property tasks, while leveraging the visual modality to directly support local and structurally complex tasks.
-To this end, we propose GraphVista, a unified framework for graph understanding with VLMs.
+Therefore, a natural conclusion is that fully leveraging VLMs for graph structure understanding requires addressing these two challenges.
+This entails, on the one hand, preserving as much task-relevant graph information as possible within limited input tokens, and on the other hand, assigning tasks according to modality strengths: letting the text modality efficiently address graph property tasks, while leveraging the visual modality to directly support local structure reasoning tasks.
+To this end, we propose GraphVista, an agent-driven unified framework designed to coordinate textual and visual modalities for robust graph understanding, ensuring the systematic robustness required to reliably tackle challenging graph tasks.
 For the scalability challenges, GraphVista is inspired by the retrieval-augmented generation (RAG) paradigm Lewis et al.
@@
 It organizes graph information hierarchically according to structural importance and stores multi-granularity textual descriptions in a lightweight GraphRAG base.
-For a given graph understanding task, GraphVista retrieves only the relevant information before input, thereby compressing redundant context and preserving key reasoning elements.
+For a given graph structure understanding task, GraphVista retrieves only the relevant information before input, thereby compressing redundant context and preserving key reasoning elements.
 When necessary, it can also generate high-resolution visual subgraphs from the retrieved local information to support complex reasoning tasks.
-To enable effective coordination across modalities, GraphVista introduces a planning agent that parses graph understanding tasks and routes them to the appropriate modality.
-The textual modality branch, enhanced with the RAG techniques, efficiently handles simple property tasks by retrieving and reasoning over concise textual descriptions.
-In contrast, the visual modality branch employs the Visual Graph Thoughts Agent, which constructs a multimodal chain of thought called Visual Graph Thoughts over high-resolution visual subgraphs, enabling VLMs to perform stepwise visual reasoning grounded in explicit structural evidence for structurally complex tasks.
-The main contributions of this paper are listed as follows:
+To enable effective coordination across modalities, GraphVista is built upon a planning agent that drives the overall workflow.
+It parses graph tasks, decomposes composite tasks if necessary, and routes them to the appropriate modality.
+The textual modality branch, enhanced with the RAG techniques, efficiently handles graph properties tasks by retrieving and reasoning over concise textual descriptions.
+In contrast, the visual modality branch employs the Visual Graph Thoughts Agent, which constructs a multimodal chain-of-thought called Visual Graph Thoughts over high-resolution visual subgraphs, enabling VLMs to perform stepwise visual reasoning grounded in explicit structural evidence for visual tasks.
+The main contributions are as follows:
   • $\bullet$
-    We present the first systematic analysis of the scalability challenges and coordination gaps that hinder VLMs in graph understanding, providing both conceptual insights and empirical evidence.
+    We present the first systematic analysis of the scalability challenges and coordination gaps that hinder VLMs in graph structure understanding, providing both conceptual insights and empirical evidence.
   • $\bullet$
-    We propose GraphVista, the first unified framework for graph understanding with VLMs, designed around two objectives: improving scalability and enabling coordination across modalities.
+    We propose GraphVista, the first unified framework for graph structure understanding with VLMs, designed around two objectives: improving scalability and enabling coordination across modalities.
 To this end, GraphVista integrates a set of innovative techniques, such as a hierarchical GraphRAG base, a planning agent for task routing, and a multimodal graph thoughts agent for complex reasoning.
   • $\bullet$
-    We introduce Grena, a scalable benchmark that supports step-level training of Visual Graph Thoughts and evaluates graph understanding from simple property to complex local tasks.
+    We present Grena, the first large-scale graph benchmark designed to support step-level evaluation of multimodal graph understanding.
   • $\bullet$
-    Extensive experiments across diverse graph understanding tasks demonstrate the effectiveness of GraphVista, highlighting its ability to scale to large graphs and to exploit the complementary strengths of textual and visual modalities.
+    Extensive experiments across diverse graph tasks demonstrate the effectiveness of GraphVista, highlighting its ability to scale to large graphs and to exploit the complementary strengths of textual and visual modalities.
 2 Related Work
-2.1 Text Modality-based Graph Understanding
-Most existing studies convert graphs into textual descriptions to enable VLM-based graph understanding, primarily focusing on benchmark construction and capability analysis.
+2.1 Text Modality-based Graph Structure Understanding
+Most existing studies convert graphs into textual descriptions to enable VLM-based graph structure understanding, primarily focusing on benchmark construction and capability analysis.
 Recent benchmarks evaluate VLMs on basic graph element comprehension Wang et al.
@@
 These methods depend on external modules rather than enhancing the intrinsic graph understanding of VLMs, and are thus orthogonal to our work.
-2.2 Visual Modality-based Graph Understanding
-Visual modality-based graph understanding is still in its infancy.
-VisionGraph Li et al.
-(2024c) and VGCure Zhu et al.
+2.2 Visual Modality-based Graph Structure Understanding
+Visual modality-based graph structure understanding is still in its infancy.
+Li et al.
+(2024c), Zhu et al.
+(2025b), and Zhao et al.
 (2025) benchmark VLMs on graph perception and reasoning, revealing difficulties in capturing structural information.
@@
 3.1 Overview
-In this section, we introduce GraphVista, an agent-driven framework designed to address the challenges of scalability and modality coordination in graph understanding, as shown in Figure 2.
-GraphVista operates under the coordination of the planning agent, which orchestrates the entire reasoning workflow.
+We introduce GraphVista, an agent-driven framework designed to address the challenges of scalability and modality coordination in graph structure understanding [^3] Aligned with prior studies Wang et al.
+(2023); Chen et al.
+(2024a); Yuan et al.
+(2025a), our work focuses on pure graph structure understanding.
+The processing of semantic information is orthogonal to our method and can be seamlessly incorporated into GraphVista due to its modular design.
+, as shown in Figure 2.
+GraphVista is coordinated by a planning agent that directs the entire workflow.
 Initially, the planning agent constructs a lightweight Hierarchical GraphRAG Base $\mathcal{K}$ by partitioning the graph based on topological centrality.
-Upon receiving a question, it parses the task and routes it to the appropriate reasoning modality.
-For Simple Property Tasks, the text modality-based branch is activated, employing GraphRAG-based retrieval to extract relevant context from $\mathcal{K}$.
-For Complex Local Tasks that require fine-grained structural reasoning, the task is delegated to the visual modality-based branch.
-The Visual Graph Thoughts Agent extracts visual subgraphs from $\mathcal{K}$ and performs iterative multimodal reasoning grounded in visual evidence.
-The rest of this section details our methodology.
-We begin by introducing the construction of $\mathcal{K}$ in §3.2.
-Next, we describe the planning agent-based task routing mechanism in §3.3.
-Finally, we elaborate on the visual modality-based and text modality-based graph understanding modules in §3.4 and §3.5, respectively.
+Upon receiving a question, the agent parses the request to identify the task type and key entities, decomposes composite tasks if necessary, and routes them to the appropriate reasoning modality.
+For text-modality tasks, the text modality-based branch is activated, employing GraphRAG-based retrieval to efficiently extract relevant context from $\mathcal{K}$.
+For visual-modality tasks, the task is delegated to the visual modality-based branch.
+In this branch, the Visual Graph Thoughts Agent extracts task-relevant visual subgraphs from $\mathcal{K}$ and performs iterative multimodal reasoning grounded in explicit visual evidence.
+The remainder of this section outlines our proposed framework.
+We begin with the construction of the hierarchical base $\mathcal{K}$ in §3.2, followed by the planning agent-driven task routing mechanism in §3.3.
+Finally, we elaborate on the graph understanding modules specific to the visual and textual modalities in §3.4 and §3.5, respectively.
 3.2 Hierarchical GraphRAG Base
-A primary challenge for LLMs and VLMs in understanding graph structures is their limited input capacity.
-This constraint makes it infeasible to encode an entire graph, often causing a loss of critical structural information.
-To address this limitation, we draw inspiration from RAG and propose building $\mathcal{K}$.
-Unlike conventional RAG pipelines Guo et al.
+Limited context windows often prevent VLMs from processing complete graphs, leading to the loss of key structural details.
+To address this, we propose constructing a local knowledge base $\mathcal{K}$.
+Diverging from conventional RAG pipelines Guo et al.
 (2025); Jin et al.
-(2025) that rely on external knowledge sources, our approach constructs a compact multi-level structural representation directly from the graph and stores it in a lightweight GraphRAG Base, enabling scalable reasoning without external dependencies [^2] For small graphs with $|V|\leq 15$, we skip storage and feed the entire graph directly into the VLMs..
-We partition the graph $G=(V,E)$ with $|V|$ nodes into three tiers based on the topological centrality of its nodes.
+(2025) that rely on external knowledge sources, our approach constructs a compact multi-level structural representation directly from the graph and stores it in a lightweight GraphRAG Base, enabling scalable reasoning without external dependencies [^4] For small graphs ($|V|\leq 15$), we skip the storage phase and feed the entire graph directly into the VLMs..
+Specifically, we partition the graph $G=(V,E)$ with $|V|$ nodes into three tiers based on the topological centrality of its nodes.
 Each tier uses a different storage granularity, significantly reducing storage needs while preserving the structural details essential for downstream tasks.
 Tier 1: Core Nodes.
-This tier consists of nodes with the highest structural influence, functioning as hubs that govern information flow and shape global graph properties.
-We employ PageRank, a widely adopted centrality measure that reflects both connectivity and reachability, to identify the top $K_{1}\%$ of nodes as the core nodes.
-For each core node $v$, its complete 2-hop neighborhood $\mathcal{N}_{2}(v)$ is preserved to capture fine-grained local structural information.
+This tier comprises nodes with the strongest structural influence, acting as hubs that govern information flow and global graph properties.
+We employ PageRank, which captures both connectivity and reachability, to identify the top $K_{1}\%$ of nodes as core nodes.
+For each, we preserve the 2-hop neighborhood [^5] We limit the storage scope to a 2-hop radius to optimize the trade-off between structural coverage and computational cost.
+Extending to 3-hops (or beyond) leads to “neighbor explosion” Barabási and Albert (1999); Grabow et al.
+(2015), yet remote nodes contribute little to local structural reasoning Li et al.
+(2018); Liben-Nowell and Kleinberg (2003).
+We adopt this generalized strategy to ensure broad applicability; advanced GraphRAG storage optimizations are orthogonal to our contribution and can be seamlessly integrated.
+$\mathcal{N}_{2}(v)$ to explicitly retain fine-grained local topology.
 Tier 2: Backbone Nodes.
-This tier contains nodes located along intercommunity paths, which are essential for preserving global connectivity.
-These nodes are selected using betweenness centrality, which quantifies how frequently a node lies on shortest paths between other nodes.
-The next $K_{2}\%$ of nodes are retained, and for each backbone node $v$, only its 1-hop neighborhood $\mathcal{N}_{1}(v)$ is stored to balance representational fidelity with storage efficiency.
+Situated along inter-community paths, these nodes are critical for maintaining global connectivity.
+We select them using Betweenness Centrality, prioritizing nodes that frequently bridge shortest paths between pairs.
+The subsequent $K_{2}\%$ of nodes are retained; however, to balance representational fidelity with storage efficiency, we store only the 1-hop neighborhood $\mathcal{N}_{1}(v)$ for each backbone node.
 Tier 3: Peripheral Nodes.
@@
 This module employs a planning agent to allocate each task to the optimal reasoning modality, either textual or visual.
-The routing procedure comprises two stages.
-In the Task Parsing stage, the agent analyzes the natural language question to extract the task intent and key entities.
-In the Task Categorization stage, the parsed task is classified and dispatched to the corresponding reasoning module.
-Task Parsing
-Given a natural language question $Q$ over $G$, the planning agent first performs semantic parsing to identify the task type $T$ and extract the set of key entities $\mathcal{E}$:
+Task Parsing.
+Given a natural language question $Q$ over $G$, the planning agent employs semantic parsing to determine the task type $T$ and extract relevant entities $\mathcal{E}$:
 $(T,\mathcal{E})=f_{\text{parse}}(Q,G)$ (1)
-, where $f_{\text{parse}}$ denotes a semantic parsing function implemented by the planning agent.
-The agent applies few-shot prompting to infer task semantics, producing a task type $T$ (e.g., neighbor identification) and a set of key entity pairs $\mathcal{E}={(v_{i},v_{j})\mid v_{i},v_{j}\in V}$ referenced in the question.
-The task type $T$ determines the routing strategy, while the entity set $\mathcal{E}$ provides essential context for downstream modules.
-To ensure uniform processing, queries involving a single entity (e.g., “What is the degree of node $v_{i}$?”) are reformulated as $(v_{i},\emptyset)$ by introducing a null entity placeholder.
-Task Categorization and Execution Planning
-Following parsing, the planning agent routes the task to the appropriate module based on $T$ and formulates a high-level execution plan [^4] The planning agent first matches the question against predefined task templates.
-If no match is found, it falls back to semantic parsing to infer the intent and choose the reasoning modality..
-Building on the task definitions established in existing graph understanding benchmarks Chen et al.
-(2024a); Yuan et al.
-(2025a); Cao et al.
-(2025), we propose a task classification framework that leverages the complementary strengths of different graph representation modalities.
-Specifically, graph tasks are categorized into two classes, each handled by a dedicated module:
-Simple Property Task Examples: “What is the total number of nodes in this Graph?”; “Return the degree of node A.” Complex Local Task Examples: “Find the shortest path between nodes A and B.”; “Do nodes A and B share any third-order neighbors?”
-Simple Property Tasks.
-These tasks involve questioning either macroscopic graph statistics or microscopic, explicitly defined properties.
-Such information is efficiently and reliably obtained from textual graph descriptions, whereas visual representations often suffer from resolution limitations that obscure fine-grained details.
-Accordingly, the planning agent routes these tasks to the GraphRAG agent module, providing a high-level retrieval strategy for execution over $\mathcal{K}$.
-Complex Local Tasks.
-These tasks require reasoning over fine-grained topological structures, such as finding the shortest path between nodes.
-For such tasks, a visual subgraph offers a concise and intuitive representation that is more effective for complex topological reasoning than verbose textual descriptions.
+, where $f_{\text{parse}}$ utilizes few-shot prompting [^7] The planning agent first matches the question against predefined task templates; if unmatched, it falls back to semantic parsing to infer the intent and choose the reasoning modality.
+Implementation details and a quantitative analysis of task classification error rates are provided in Appendix H.
+to infer task semantics, producing a task category $T$ (e.g., neighbor identification) and a set of entity pairs $\mathcal{E}$.
+To ensure consistent downstream processing, queries involving a single entity (e.g., node degree) are normalized to $(v_{i},\emptyset)$ using a null placeholder.
+Task Routing and Execution Planning.
+After parsing the task, the planning agent routes the problem to the appropriate processing module based on the task type $T$ and formulates a detailed execution plan.
+Integrating task designs from existing graph understanding benchmarks, the planning agent routes tasks based on their dependence on structural reasoning, thereby aligning each task with the respective strengths of text and vision modalities [^8] Composite tasks that can be decomposed into single-modality operations (e.g., determining which of two nodes shares more common neighbors with a given node) can be handled by the planning agent through sequential execution of the corresponding subtasks..
+Specifically, tasks that do not require structural reasoning are handled by the text modality, tasks requiring local structural reasoning are assigned to the visual modality, and tasks involving global graph structure are decomposed into stepwise text- and vision-based subtasks.
+Text-Modality Task Examples: “What is the total number of nodes in this graph?”; “What is the degree of node 15?” Visual-Modality Task Examples: “Find the shortest path between nodes A and B.”; “Whether node A is part of any triangle.” Modality-Collaborative Task Examples: “Calculate the diameter of the graph.” Decomposition: (i) Identify peripheral nodes using text; (ii) Estimate pairwise distances through visual subgraph reasoning.”
+Text-Modality Tasks.
+These tasks do not require structural reasoning over the graph.
+Instead, they rely on direct access to individual nodes or edges for basic statistical retrieval, such as counting the total number of nodes or querying the degree of a given node.
+Accordingly, the planning agent directs these tasks to the GraphRAG agent module, providing a retrieval strategy for execution over $\mathcal{K}$.
+Visual-Modality Tasks.
+These tasks involve reasoning over local graph structures rather than simple retrieval.
+They focus on topological relationships within a limited subgraph (e.g., shortest-path tasks).
+In such cases, visual subgraph representations provide a clearer and more efficient form for structural reasoning than textual descriptions.
 Consequently, the planning agent delegates such tasks to the visual graph thoughts agent, while decomposing the task into algorithm-guided substeps and constructing a high-level reasoning plan.
-3.4 Visual Modality-based Graph Understanding
-This module leverages visual-modality graph representations to handle complex local problems that require fine-grained topological understanding and multi-step logical inference.
-To address the scalability issue of large graphs, we extract a task-relevant $k$-hop subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with a maximum of $N_{max}$ nodes from $G$ in $\mathcal{K}$, centered on the key entities $\mathcal{E}$ (details are provided in Appendix B).
-The resulting subgraph is then transformed into a high-resolution visual representation $G_{\text{image}}=f_{\mathrm{viz}}(G^{\prime})$, which preserves the essential topological structure of $G^{\prime}$ and serves as a faithful basis for downstream reasoning.
+Modality-Collaborative Tasks.
+These tasks rely on global graph structure understanding that combines full-graph access with fine-grained structural analysis, which is often beyond a single modality.
+They are therefore formulated as a sequence of steps that alternate between global information access and local structural reasoning.
+The planning agent decomposes the task into sequential text- and visual-modality sub-tasks.
+3.4 Visual Modality-based Graph Structure Understanding
+This module leverages visual-modality graph representations to handle visual-modality tasks that require fine-grained topological understanding and multi-step logical inference.
+To address scalability, we extract a task-relevant $k$-hop subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ centered on key entities $\mathcal{E}$, with at most $N_{\max}$ nodes (as detailed in Appendices B and C).
+$G^{\prime}$ is visualized as $G_{\text{image}}=f_{\mathrm{viz}}(G^{\prime})$ via a task-driven strategy that adapts layouts to the task and simplifies dense subgraphs for visual clarity, ensuring that essential structural information remains perceivable to the VLM.
 3.4.1 Visual Graph Thoughts
@@
 (2025).
-To overcome this limitation in multimodal reasoning for complex local tasks, we introduce Visual Graph Thoughts.
+To overcome this limitation in multimodal reasoning, we introduce Visual Graph Thoughts.
 After obtaining the visual subgraph $G_{\text{image}}$, the visual reasoning agent $\mathcal{M}_{\text{VRA}}$, instantiated from a VLM, performs iterative multimodal reasoning under the guidance of a high-level plan $\Pi=(\pi_{1},\pi_{2},\dots,\pi_{n})$ generated by the planning agent.
@@
 Conditioned on $S_{t}$, $\mathcal{M}_{\text{VRA}}$ produces an intermediate reasoning output $o_{t}$ and an action $a_{t}$:
-$(o_{t},a_{t})=\mathcal{M}_{\text{VRA}}(S_{t})=\mathcal{M}_{\text{VRA}}(G_{\text{image}}^{(t-1)},H_{t-1},{\pi}_{t})$ (2)
+$(o_{t},a_{t})=\mathcal{M}_{\text{VRA}}(S_{t})=\mathcal{M}_{\text{VRA}}(G_{\text{image}}^{(t-1)},H_{t-1},{\pi}_{t}).$ (2)
 If $a_{t}$ invokes $f_{\mathrm{viz}}$ (e.g., node highlighting), the visual representation is updated as $G_{\text{image}}^{(t)}=f_{\mathrm{viz}}(G_{\text{image}}^{(t-1)},a_{t})$; otherwise, $G_{\text{image}}^{(t)}=G_{\text{image}}^{(t-1)}$.
 This closed-loop process continues until all plan steps are executed.
-Finally, $\mathcal{M}_{\text{VRA}}$ aggregates the reasoning trace $H_{n}$ to produce the final answer $A_{\text{final}}=f_{\text{summarize}}(H_{n})$, where $f_{\text{summarize}}$ aggregates intermediate reasoning outputs to generate the final prediction.
+Finally, $\mathcal{M}_{\text{VRA}}$ aggregates the reasoning trace $H_{n}$ to produce the final answer.
 3.4.2 Aligning Visual Graph Thoughts with Process-level DPO
-To further improve the multimodal reasoning capability of $\mathcal{M}_{\text{VRA}}$ by promoting the generation of reliable visual graph thoughts, we adopt a process-level Direct Preference Optimization (DPO) training strategy Rafailov et al.
-(2023).
-Compared to outcome-based methods Kong et al.
-(2025a); Liu et al.
-(2025), we model the entire multimodal reasoning trajectory, which provides stable, step-level reinforcement for graph inference tasks.
+To further improve the multimodal reasoning capability of $\mathcal{M}_{\text{VRA}}$ by promoting the generation of reliable visual graph thoughts, we adopt the process-level Direct Preference Optimization (DPO) training strategy Rafailov et al.
+(2023), which models the full multimodal reasoning trajectory and provides stable, step-level supervision for graph inference.
+Advanced reinforcement learning methods Hu et al.
+(2025); Dinucu-Jianu et al.
+(2025); Yu et al.
+(2025) are orthogonal to our framework and can be naturally integrated.
+We adopt the standard process-level DPO formulation to ensure generality and training stability.
 1. Construction of Process Preference Dataset.
 We construct a process preference dataset $\mathcal{D}$, where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$.
-Here, $y_{w}$ denotes the “Chosen” path representing the preferred reasoning trajectory, while $y_{l}$ denotes the “Rejected” path representing a non-preferred trajectory.
+Here, $\mathbf{x}$ denotes the multimodal input, $y_{w}$ represents the “Chosen” path (preferred reasoning trajectory), and $y_{l}$ represents the “Rejected” path (non-preferred trajectory).
 The construction details are in Appendix A.
@@
 , where the policy ratio is defined as $r_{\theta}(y|\mathbf{x})=\pi_{\theta}(y|\mathbf{x})/\pi_{\mathrm{ref}}(y|\mathbf{x})$, and $\beta$ controls the regularization strength.
-By maximizing the log-probability margin between the preferred ($y_{w}$) and rejected ($y_{l}$) trajectories, the model is guided away from predefined reasoning errors.
-The resulting policy generates visual graph thoughts that are both logically coherent and tightly grounded in visual graph evidence throughout the inference process.
-3.5 Text Modality-based Graph Understanding
-For simple property tasks, we adopt GraphRAG-based retrieval for targeted information extraction.
+Maximizing the log-probability margin between preferred ($y_{w}$) and rejected ($y_{l}$) trajectories effectively suppresses predefined reasoning errors.
+The resulting policy generates visual graph thoughts that are logically coherent and faithfully grounded in visual evidence.
+Training details are available in Appendices E and F.
+3.5 Text Modality-based Graph Structure Understanding
+For text-modality tasks, we adopt GraphRAG-based retrieval for targeted information extraction.
 Note that existing RAG variants and optimizations Deng et al.
@@
 To maintain generality, we employ the standard RAG formulation in this work.
-For a given task request $(Q,\mathcal{E})$, we first employs a retrieval function $f_{\text{retrieve}}$ to question $\mathcal{K}$ and extract relevant context, formalized as $C_{\text{retrieved}}=f_{\text{retrieve}}(\mathcal{K},(Q,\mathcal{E}))$.
-Subsequently, the downstream VLM is prompted with both the original question $Q$ and the retrieved context $C_{\text{retrieved}}$ to generate the final answer.
-Table 1: Performance comparison on Grena and GraphSQA Benchmark. Red and Blue highlight the best and second-best results respectively.
+Given the parsed tuple $(Q,\mathcal{E})$, the relevant context $\mathcal{C}$ is extracted from $\mathcal{K}$ by aggregating the structural descriptions of entities in $\mathcal{E}$.
+The VLM then generates the answer conditioned on both $Q$ and $\mathcal{C}$.
+4 Evaluation
+Table 1: Performance comparison on Grena and GraphSQA Benchmark. Red and Blue highlight the best and second-best results, respectively.
 Method Model Grena GraphSQA
-Simple Complex Overall Simple Complex Overall
-Text-only InternVL3-9B 0.1211 0.0287 0.0829 0.4788 0.2498 0.3169
-GLM-4.1V-9B 0.0157 0.0071 0.0121 0.2532 0.1670 0.1929
-Qwen2.5-VL-7B 0.0096 0.0095 0.0096 0.2340 0.1122 0.1488
-Qwen3-8B 0.0065 0.0056 0.0061 0.1351 0.1352 0.1352
-Gemma-3-12B 0.0369 0.0358 0.0364 0.5917 0.3245 0.4033
-GraphPRM (DPO) Qwen2.5-VL-7B (DPO) 0.0332 0.0087 0.0231 0.7198 0.3455 0.4550
-Qwen3-8B (DPO) 0.0060 0.0049 0.0055 0.3629 0.2357 0.2728
-Gemma-3-12B (DPO) 0.0417 0.0335 0.0383 0.6235 0.3455 0.4195
-GraphInsight InternVL3-9B 0.1788 0.0056 0.1073 0.7564 0.2315 0.3860
-GLM-4.1V-9B 0.1849 0.0016 0.1092 0.4313 0.3273 0.3583
-Qwen2.5-VL-7B 0.1925 0.0085 0.1165 0.5769 0.2293 0.3348
-Gemma-3-12B 0.2919 0.0236 0.1811 0.7372 0.2478 0.3779
-GraphToken Qwen2.5-VL-7B 0.0981 0.0782 0.0899 0.0388 0.1770 0.1376
-Qwen3-8B 0.1589 0.0757 0.1245 0.1742 0.1508 0.1579
-Gemma-3-12B 0.0158 0.0576 0.0331 0.0566 0.1091 0.0927
-GITA InternVL3-9B 0.1751 0.1098 0.1481 0.3910 0.1755 0.2402
-GLM-4.1V-9B 0.1651 0.1175 0.1454 0.3397 0.1562 0.2112
-Qwen2.5-VL-7B 0.1351 0.0632 0.1054 0.3205 0.1596 0.2079
-Gemma-3-12B 0.1732 0.1435 0.1609 0.2692 0.1739 0.2026
-GraphVista (Ours) InternVL3-9B 0.9357 0.3578 0.6970 0.7970 0.6886 0.7211
-GLM-4.1V-9B 0.9387 0.2727 0.6637 0.7954 0.6657 0.7046
-Gemma-3-12B 0.9362 0.3477 0.6932 0.7991 0.6815 0.7168
-Gemma-3-4B (DPO) 0.9362 0.3397 0.6899 0.7943 0.6780 0.7129
-Qwen2.5-VL-7B (DPO) 0.9367 0.3912 0.7114 0.7973 0.6806 0.7156
-4 Evaluation
+Text Visual Collab. Overall Text Visual Overall
+Text-only GLM-4.1V-9B 0.0352 0.0055 0.0110 0.0217 0.2532 0.1670 0.1929
+InternVL3-9B 0.0452 0.0180 0.0303 0.0345 0.4788 0.2498 0.3185
+Qwen2.5-VL-7B 0.0558 0.0345 0.0337 0.0450 0.2340 0.1122 0.1487
+Qwen3-8B 0.0403 0.0367 0.0096 0.0321 0.1351 0.1352 0.1352
+Gemma-3-12B 0.1246 0.0591 0.0337 0.0858 0.5917 0.3245 0.4047
+GPT-5-mini 0.1451 0.0933 0.0515 0.1093 0.6464 0.6082 0.6197
+GraphPRM (DPO) Qwen2.5-VL-7B (DPO) 0.0589 0.0410 0.0314 0.0477 0.7198 0.3455 0.4578
+Qwen3-8B (DPO) 0.0446 0.0379 0.0068 0.0339 0.3629 0.2357 0.2739
+Gemma-3-12B (DPO) 0.1281 0.0590 0.0434 0.0899 0.6235 0.3455 0.4289
+GraphInsight GLM-4.1V-9B 0.1951 0.2655 0.2172 0.2189 0.4313 0.3273 0.3585
+InternVL3-9B 0.2507 0.2915 0.2088 0.2515 0.7564 0.2315 0.3890
+Qwen2.5-VL-7B 0.2011 0.1078 0.0976 0.1520 0.5769 0.2293 0.3336
+Gemma-3-12B 0.2469 0.2357 0.1131 0.2123 0.7372 0.2478 0.3946
+GPT-5-mini 0.3031 0.3429 0.2843 0.3091 0.7211 0.6446 0.6676
+GraphToken Qwen2.5-VL-7B 0.0141 0.0982 0.0396 0.0423 0.0388 0.1770 0.1355
+Qwen3-8B 0.0052 0.0665 0.0089 0.0222 0.1742 0.1508 0.1578
+Gemma-3-12B 0.0191 0.1064 0.0475 0.0488 0.0566 0.1091 0.0933
+GITA GLM-4.1V-9B 0.0665 0.2152 0.1835 0.1333 0.3397 0.1562 0.2112
+InternVL3-9B 0.0714 0.1045 0.1546 0.0998 0.3910 0.1755 0.2401
+Qwen2.5-VL-7B 0.1099 0.1799 0.2271 0.1561 0.3205 0.1596 0.2079
+Gemma-3-12B 0.1218 0.1591 0.2344 0.1583 0.2692 0.1739 0.2025
+GraphVista (Ours) GLM-4.1V-9B 0.9352 0.3540 0.2559 0.6214 0.7954 0.6657 0.7046
+InternVL3-9B 0.9357 0.3701 0.3449 0.6469 0.7970 0.6886 0.7211
+Qwen2.5-VL-7B (DPO) 0.9363 0.4309 0.4478 0.6876 0.7973 0.6806 0.7156
+Gemma-3-12B (DPO) 0.9361 0.4431 0.6018 0.7272 0.7991 0.6815 0.7168
+GPT-5-mini 0.9406 0.4526 0.6241 0.7372 0.8497 0.7967 0.8126
 4.1 Experimental Setup
-4.1.1 Baselines
-We compare GraphVista with representative state-of-the-art baselines across three categories: (a) Text-based methods, which reason solely over textual graph descriptions, including the standard Text-only approach, GraphPRM Peng et al.
-(2025) trained via DPO, and GraphInsight Cao et al.
+Baselines.
+We compare GraphVista with state-of-the-art baselines across three categories: (a) Text-based methods, which reason exclusively over textual descriptions, including the standard Text-only baseline, GraphPRM Peng et al.
+(2025) (trained via DPO), and GraphInsight Cao et al.
 (2025).
-(b) GNN-based methods, which encode structural information into latent representations using graph neural networks.
-A representative example is GraphToken [^5] These methods are orthogonal to GraphVista, focusing on graph encoding rather than improving the intrinsic graph understanding of VLMs.
-We include them for completeness.
-Perozzi et al.
-(2024).
-(c) Hybrid methods, exemplified by GITA Wei et al.
-(2024), which integrate both visual and textual graph modalities into VLMs.
-4.1.2 Benchmarks and Evaluation Tasks
-Existing graph understanding benchmarks suffer from limited coverage of graph scales and topologies, a lack of task categorization grounded in fundamental graph properties, and insufficient support for evaluating multi-step visual reasoning.
-To address these gaps, we introduce Grena, a large-scale, multi-task benchmark we designed to evaluate graph understanding of LLMs and VLMs, which encompasses a wide range of task types found in existing benchmarks Wang et al.
+(b) GNN-based methods, such as GraphToken Perozzi et al.
+(2024), which encode structural information into latent representations [^9] These methods focus on graph encoding rather than VLM intrinsic understanding.
+We include them for completeness..
+(c) Hybrid methods, represented by GITA Wei et al.
+(2024), which integrate both visual and textual modalities.
+Benchmarks and Evaluation Metrics.
+Current graph understanding benchmarks suffer from restricted scale and topological diversity, with insufficient support for multi-step visual reasoning.
+We introduce Grena, a large-scale benchmark designed to evaluate VLMs across 20 distinct task types, covering the scope of existing baselines Wang et al.
 (2023); Yuan et al.
 (2025a); Zhu et al.
-(2025).
-Grena spans a wide diversity of graph scales, with node counts ranging up to 2050, and encompasses 18 distinct task types, which are organized into two categories: (1) 7 Simple Property Tasks (Simple) focus on querying basic properties and local structures, and (2) 11 Complex Local Tasks (Complex) require fine-grained topological information and performing fine-grained inference.
+(2025b).
+Grena spans a wide range of graph scales (up to 2,050 nodes) and topologies.
 Further details are provided in Appendix A.
-In our experiments, we use Grena to evaluate graph understanding performance on large-scale graphs ($|V|\in[50,2050]$), covering a total of 13,860 tasks, which are split into 8,136 simple and 5,724 complex instances.
+We use Grena to evaluate graph structure understanding performance on large-scale graphs ($|V|\in[50,2050]$), covering a total of 22,800 tasks.
 Complementarily, we use the GraphSQA benchmark Cao et al.
-(2025) for small-graph understanding, with $|V|$ in the range of $[15,50]$.
+(2025) for small-graph structure understanding, with $|V|$ in the range of [15,50][15,50] [^10] As this small-graph benchmark lacks modality-collaborative tasks, we do not separate local and global tasks.
+Instead, we directly adopt its original classification, referring to the two categories as “Text” and “Visual”..
 Across all tasks and benchmarks, we use Accuracy (ACC) as the evaluation metric.
-Methods that require DPO are trained on $\mathcal{D}$.
-(a) Complex (Gemma-3)
-(b) Simple (Gemma-3)
-(c) Complex (Qwen2.5-VL)
-(d) Simple (Qwen2.5-VL)
-Figure 3: Analysis on Graphs with Different $|V|$
-Table 2: Ablation study on reasoning strategies of GraphVista. The best results within each VLM group are bolded, and the second-best results are underlined.
-Strategy Simple Complex Overall
+Methods requiring DPO are trained on $\mathcal{D}$, as detailed in Appendix E.
+Table 2: Ablation study of reasoning strategies in GraphVista. Best results within each VLM group are bolded, and second-best results are underlined.
+Strategy Text Visual Collab. Overall
+GLM-4.1V-9B
+CoT (Baseline) 0.9352 0.1376 0.1667 0.5433
+Visual Graph Thoughts (Text) 0.2734 0.1902 0.5846
+Visual Graph Thoughts (Visual) 0.3540 0.2559 0.6214
 InternVL3-9B
-CoT (Baseline) 0.9368 0.1422 0.6086
-Visual Graph Thoughts (Text) 0.2220 0.6416
-Visual Graph Thoughts (Visual) 0.3578 0.6977
-Gemma-3-12B
-CoT (Baseline) 0.9362 0.1406 0.6076
-Visual Graph Thoughts (Text) 0.2447 0.6506
-Visual Graph Thoughts (Visual) 0.3477 0.6932
-GLM-4.1V-9B
-CoT (Baseline) 0.9387 0.1406 0.6091
-Visual Graph Thoughts (Text) 0.2300 0.6460
-Visual Graph Thoughts (Visual) 0.2727 0.6637
+CoT (Baseline) 0.9357 0.1422 0.1332 0.5368
+Visual Graph Thoughts (Text) 0.2879 0.2511 0.6031
+Visual Graph Thoughts (Visual) 0.3701 0.3449 0.6469
 Qwen2.5-VL-7B (DPO)
-CoT (Baseline) 0.9367 0.1414 0.6083
-Visual Graph Thoughts (Text) 0.3429 0.6915
-Visual Graph Thoughts (Visual) 0.3912 0.7114
-4.1.3 Models
-We instantiate GraphVista and all baselines using recent open-source, competitive VLMs, including Gemma-3-12B, Gemma-3-4B Kamath et al.
+CoT (Baseline) 0.9363 0.1691 0.1684 0.5525
+Visual Graph Thoughts (Text) 0.3629 0.3355 0.6431
+Visual Graph Thoughts (Visual) 0.4309 0.4478 0.6876
+Gemma-3-12B (DPO)
+CoT (Baseline) 0.9361 0.1676 0.1703 0.5525
+Visual Graph Thoughts (Text) 0.3551 0.4200 0.6610
+Visual Graph Thoughts (Visual) 0.4431 0.6018 0.7272
+Models.
+We instantiate all methods using recent open-source, competitive VLMs, including Gemma-3-12B Kamath et al.
 (2025), Qwen2.5-VL-7B Bai et al.
-(2025), InternVL3-9B Zhu and et al (2025), and GLM-4.1V-9B-Thinking Hong et al.
+(2025), InternVL3-9B Zhu et al.
+(2025a), and GLM-4.1V-9B-Thinking Hong et al.
 (2025), a reasoning-oriented VLM.
@@
 (2025) to ensure a fairer comparison.
+For training-free methods, we include GPT-5-mini to assess generality.
+Table 3: Ablation comparing frozen base models (“Frozen”) and DPO-fine-tuned models (“DPO”). Red and Blue denote the best and second-best results.
+Model Strategy Text Visual Collab. Overall
+Qwen2.5-VL-7B Frozen 0.9363 0.3951 0.3589 0.6571
+DPO 0.4309 0.4478 0.6876
+Gemma-3-12B Frozen 0.9361 0.4005 0.5387 0.7010
+DPO 0.4431 0.6018 0.7272
+(a) Text (Gemma-3)
+(b) Text (Qwen2.5-VL)
+(c) Visual (Gemma-3)
+(d) Visual (Qwen2.5-VL)
+(e) Collab. (Gemma-3)
+(f) Collab. (Qwen2.5-VL)
+Figure 3: Analysis on Graphs with Different $|V|$
 4.2 Performance
-As shown in Table 1, GraphVista significantly outperforms all baseline methods on both Grena and GraphSQA benchmarks.
-4.2.1 Simple Property Tasks
+As shown in Table 1, GraphVista consistently outperforms all baselines on both benchmarks.
+Text-modality Tasks.
 […]
-4.2.2 Complex Local Tasks
+Visual and Modality-Collaborative Tasks.
 […]
-Table 3: Ablation results comparing frozen base models (“Frozen”) with models further optimized via DPO (“DPO”).Red and Blue highlight the best and second-best results respectively.
-Model Strategy Simple Complex Overall
-Gemma-3-4B Frozen 0.9362 0.2933 0.6707
-DPO 0.3397 0.6899
-Qwen2.5-VL-7B Frozen 0.9367 0.3240 0.6837
-DPO 0.3912 0.7114
 4.3 Ablation Study
-4.3.1 Effectiveness of Reasoning Strategies
-We evaluate the effect of visual grounding by comparing Visual Graph Thoughts (Visual) with its text-only variant and the standard textual CoT, where visual subgraphs are replaced by textual descriptions.
-In Table 2, the visual variant consistently outperforms the baselines, especially on complex tasks, achieving up to $1.1\times$ and $2.8\times$ higher accuracy than the text-only and CoT settings, respectively.
-These highlight the importance of explicit structural grounding, as purely textual reasoning often yields ambiguous interpretations.
-4.3.2 Impact of DPO
-We evaluate the impact of process-level DPO by comparing fine-tuned models with their frozen-base counterparts.
-As reported in Table 3, DPO enhances performance across all settings, with the most significant gains on complex tasks.
-For instance, DPO boosts the accuracy of Qwen2.5-VL-7B on complex tasks from 0.3240 to 0.3912.
-These results validate that aligning model generation with expert visual reasoning trajectories is highly effective for complex local tasks.
-(a) Gemma-3 (b) Qwen2.5-VL
-Figure 4: Hyperparameter analysis on $k$ and $N_{max}$
-Figure 5: Hyperparameter analysis on storage strategy of $\mathcal{K}$
+Reasoning Strategies.
+[…]
+Impact of DPO.
+Comparing fine-tuned models with frozen baselines (Table 3) demonstrates that process-level DPO enhances performance across all settings, particularly for complex tasks.
+Notably, DPO improves Qwen2.5-VL-7B’s accuracy on collaborative tasks from 0.3589 to 0.4478.
+This verifies the effectiveness of aligning model outputs with expert visual reasoning trajectories.
+(a) Gemma-3 (Visual)
+(b) Gemma-3 (Collab.)
+(c) Qwen2.5-VL (Visual)
+(d) Qwen2.5-VL (Collab.)
+Figure 4: Hyperparameter analysis on $k$ and $N_{\max}$
+Figure 5: Tier number analysis in $\mathcal{K}$
 4.4 Hyperparameter Analysis
-4.4.1 Subgraph Extraction Parameters $k$ and $N_{max}$
-We analyze the impact of $k$ and $N_{max}$ on performance for complex local tasks.
+Subgraph Extraction.
+We analyze the impact of $k$ and $N_{max}$ on performance for visual-modality tasks.
 As shown in Figure 4, Gemma-3 reaches its highest accuracy with a 2-hop neighborhood, indicating that it can leverage a broader structural context.
@@
 These observations point to the need for adapting subgraph extraction strategies to the characteristics of different VLM architectures.
-4.4.2 Storage Strategy of $\mathcal{K}$
-We analyze the effect of pruning node tiers in $\mathcal{K}$, which are defined by the parameters $K_{1}$% and $K_{2}$%.
-As demonstrated in Figure 5, performance on simple property tasks improves as more tiers of the graph are stored.
-Retaining just the top $K_{1}$% of nodes (Tier 1) along with their 2-hop neighborhoods already provides a strong baseline for localized retrieval.
-Adding the next $K_{2}$% of nodes (Tier 2) enhances performance further, although the gains diminish when including peripheral nodes (Tier 3).
-While full graph storage is essential for precise macroscopic tasks where any pruning leads to information loss, these findings reveal a clear trade-off between performance and storage efficiency for general reasoning tasks.
+Number of Tiers in $\mathcal{K}$.
+We examine the impact of storage granularity in $\mathcal{K}$ across three settings: Tier 1 (Core), Tier 1+2 (Backbone), and Full Graph.
+As shown in Figure 5, performance positively scales with structural coverage.
+Tiers 1 and 2 provide efficient baselines, but adding Tier 3 significantly improves performance (e.g., 0.936 on InternVL3-9B) at the cost of higher storage.
+This underscores a clear trade-off: prioritizing central nodes ensures efficiency, whereas full graph coverage is essential for maximizing accuracy.
 5 Conclusion
-In this paper, we propose GraphVista, a unified framework addressing scalability and modality coordination in graph understanding.
-To improve scalability, GraphVista uses a lightweight hierarchical GraphRAG base to compress and store task-relevant information.
-For modality coordination, the planning agent routes tasks to the most suitable modality, text modality for simple property tasks or visual modality for complex local tasks.
-Future work will extend GraphVista to more complex graph types with labels and semantics.
+In this paper, we propose GraphVista, a unified framework for scalable and modality-coordinated understanding of graph structure.
+For scalability, it employs a hierarchical GraphRAG base to compress task-relevant information.
+For coordination, a planning agent routes tasks to the optimal modality, leveraging textual retrieval for properties and visual graph thoughts for topological reasoning.
+Future work will extend GraphVista to complex graphs with labels and semantics.
 Limitations
-The proposed GraphVista framework offers a promising approach for scalable, modality-coordinated graph understanding, while also presenting several potential extensions that could inspire future research directions.
+The proposed GraphVista framework offers a promising approach for scalable, modality-coordinated graph structure understanding, while also presenting several potential extensions that could inspire future research directions.
 First, while the framework currently focuses on graph topology, its modular design is intended to facilitate extensions for handling rich semantic information.
@@
 Appendix A Grena Benchmark
-Existing graph understanding benchmarks suffer from three major limitations: (i) limited coverage of graph scales and topological structures; (ii) lack of task categorization grounded in fundamental graph properties; and (iii) insufficient support for modeling and evaluating multi-step visual reasoning in graph inference.
+Existing graph structure understanding benchmarks suffer from three major limitations: (i) limited coverage of graph scales and topological structures; (ii) lack of task categorization grounded in fundamental graph properties and reasoning modalities; and (iii) insufficient support for modeling and evaluating multi-step visual reasoning in graph inference.
 To address these gaps, we propose the Grena Benchmark, a large-scale, multi-task benchmark with step-level visual reasoning capability.
-It is designed to rigorously evaluate and advance the cognitive and reasoning abilities of Vision-language models (VLMs) in graph domains.
+It is designed to rigorously evaluate and advance the cognitive and reasoning abilities of VLMs in graph domains.
 The core design of Grena Benchmark is reflected in the following three aspects:
@@
 A.1 Scalability and Structural Diversity
-A primary objective of Grena is to systematically evaluate the scalability graph understanding capabilities of VLMs.
+A primary objective of Grena is to systematically evaluate the scalability graph structure understanding capabilities of VLMs.
 To this end, the benchmark includes graphs that span a wide range of scales, with node counts from 50 to 2,050.
@@
 This combination of scales and topologies creates a robust testbed for evaluating model generalization across heterogeneous graph structures.
-A.2 Categorization of Graph Understanding Tasks
-To provide a structured assessment of reasoning abilities, Grena Benchmark organizes tasks into two categories:
-  • Simple QA Tasks.
-These tasks focus on querying basic properties and local structures of graphs, such as “What is the total number of nodes?” or “What is the degree of node $v_{A}$?” They primarily evaluate a model’s ability to store and retrieve graph information.
-With a highly regularized form, these tasks emphasize breadth of coverage and scalability.
-  • Complex Reasoning Tasks.
-These tasks require models to integrate fine-grained topological structures and perform logical reasoning.
-For example, shortest-path problems demand identifying the optimal path from node $v_{A}$ to node $v_{B}$.
-Compared with retrieval tasks, these tasks probe deeper reasoning capabilities by requiring structural integration and inference.
+A.2 Categorization of Graph Structure Understanding Tasks
+To provide a structured assessment of reasoning abilities across different modalities, Grena Benchmark organizes tasks into three categories:
+  • Text-Modality Tasks.
+These tasks focus on basic statistical retrieval and accessing individual node attributes without requiring reasoning over complex topology.
+Examples include “What is the total number of nodes?” or “What is the degree of node $v_{A}$?” They primarily evaluate a model’s ability to interact with the Knowledge Graph ($\mathcal{K}$) for direct information retrieval.
+  • Visual-Modality Tasks.
+These tasks involve reasoning over local graph structures and topological relationships within a limited subgraph.
+Examples include shortest-path problems (“Find the shortest path between nodes A and B”) or cycle detection (“Is node A part of any triangle?”).
+Unlike retrieval tasks, these require the model to utilize visual subgraph representations for efficient structural inference.
+  • Modality-Collaborative Tasks.
+These tasks require integrating global graph understanding with fine-grained local structural analysis, necessitating a sequence of steps that alternate between modalities.
+For example, calculating the graph diameter involves identifying peripheral nodes via text modality followed by estimating pairwise distances via visual modality.
+These tasks evaluate the model’s ability to decompose complex queries into sequential sub-tasks.
 A.3 Process-Level Supervision for Visual Reasoning
@@
 The resulting paths feature common reasoning pitfalls (as detailed in Table 5), teaching the model to distinguish between correct and flawed reasoning.
-Table 6: Task types in the Grena Benchmark: definition, expected output and category.
+Table 6: Taxonomy of tasks in the Grena Benchmark, covering 20 distinct graph structure understanding challenges.
 Task Name Definition and Expected Output
-Bipartite Detection Determine whether $G=(V,E)$ is bipartite (2-colorable). Output: boolean (is bipartite?), optionally a bipartition $(S,T)$.
-Clique Detection Detect the existence of (or find) a clique of specified size $k$, or return maximal cliques. Output: boolean (exists $k$-clique?) or node set(s) forming a clique.
-Common Third-Order Neighbors Given nodes $u,v$, compute the intersection $N^{(3)}(u)\cap N^{(3)}(v)$ where $N^{(3)}(\cdot)$ denotes nodes at distance exactly 3. Output: set of nodes.
-Connectivity Check Decide whether $G$ is connected (single connected component). Output: boolean.
-Connectivity Analysis Full analysis of connectivity: number of connected components, sizes of components, and component ID/size for a query node.
-Output: integer(s) and list(s) (component counts and sizes).
-Critical Node Detection Identify articulation points (nodes whose removal increases the number of connected components) or nodes critical to connectivity. Output: set of node ids.
-Cycle Detection Determine whether any cycle exists; optionally return an example cycle (sequence of nodes). Output: boolean and, if requested, a cycle path.
-Edge Count Return the total number of edges $|E|$. Output: integer.
-Edge Existence Given a pair $(u,v)$, decide whether edge $(u,v)\in E$. Output: boolean.
-Find Connected Edges Given a node $v$ (or a small subgraph), return the set of incident edges or the edges that connect the queried substructure. Output: set of edges (pairs).
-Highest-Degree Neighbor For a query node $v$, return the neighbor $w\in N(v)$ with the largest degree $\deg(w)$ (tie-breaking rule: smallest id or deterministic rule). Output: node id.
-Neighbor Connections Measure connections among the neighbors of $v$: e.g., number of edges in the induced subgraph $G[N(v)]$. Output: integer (edge count) or list of edges.
-Node Count Return the total number of nodes $|V|$. Output: integer.
-Node Degree Given node $v$, return its degree $\deg(v)$ (or weighted degree if weights provided). Output: integer (or numeric).
-Planarity Testing Determine whether $G$ is planar; optionally, produce a planar embedding if planar. Output: boolean and (optional) embedding.
-Shortest Path Given nodes $u,v$, compute a shortest path (and its length). Supports unweighted or weighted variants. Output: sequence of nodes (path) and length.
-Third-Order Neighbors For node $v$, return $N^{(3)}(v)$, the set of nodes at graph distance exactly 3 from $v$. Output: set of nodes.
-Triangle Counting Count all triangles (3-cycles) in $G$, and/or enumerate them. Output: integer (triangle count) or list of 3-tuples (node triples).
+Articulation Point Detection Identify nodes whose removal increases the number of connected components. Output: List of node IDs.
+Common 3rd-Order Neighbor ID Find nodes that are exactly 3 hops away from both query nodes $u$ and $v$. Output: Set of node IDs.
+Connectivity Detection Determine whether the graph is fully connected. Output: Boolean.
+Cycle Detection Determine whether the graph contains any cycles. Output: Boolean.
+Edge Counting Return the total number of edges $|E|$. Output: Integer.
+Edge Existence Checking Determine if a direct edge exists between a given pair of nodes $(u,v)$. Output: Boolean.
+Connected Edge Identification Return all edges incident to a specific query node $v$. Output: Set of edges.
+Graph Diameter Calculation Compute the longest shortest path between any two nodes in the graph. Output: Integer.
+Highest Degree Neighbor ID For a query node $v$, identify the neighbor with the maximum degree. Output: Node ID.
+Hub Node Path Finding Identify a path between two nodes that routes through high-degree hub nodes. Output: Node sequence.
+Maximum Degree Node ID Identify the node(s) with the highest degree in the entire graph. Output: Node ID(s).
+Maximum Clique Detection Find the size or set of nodes of the largest complete subgraph. Output: Integer (size) or set of nodes.
+Neighbor Connection Analysis Count the number of edges within the induced subgraph of a node’s neighbors. Output: Integer.
+Node Counting Return the total number of nodes $|V|$. Output: Integer.
+Node Degree Identification Return the degree of a specific query node $v$. Output: Integer.
+Planarity Testing Determine if the graph can be embedded in a plane without edge crossings. Output: Boolean.
+Shortest Path Finding Compute the shortest path distance and sequence between nodes $u$ and $v$. Output: Path length and sequence.
+Star Structure Identification Determine if a specific subgraph forms a star topology centered at node $v$. Output: Boolean.
+Third-Order Neighbor ID Return the set of nodes at exactly distance 3 from a query node $v$. Output: Set of node IDs.
+Triangle Counting Count the total number of triangles (3-cycles) in the graph. Output: Integer.
 Appendix B Subgraph Extraction
-To mitigate the scalability challenges of large graphs, we first extract a task-relevant local subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ from $G$ stored in $\mathcal{K}$, based on the key entities $\mathcal{E}$, formalized as:
+To mitigate the scalability challenges of large graphs, we first extract a task-relevant local subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ from $G$ stored in $\mathcal{K}$, based on the key entities $\mathcal{E}$.
+This process is formalized as:
 $G^{\prime}=f_{\text{extract}}(G,\mathcal{E})$ (4)
-, where the extraction strategy $f_{\text{extract}}$ is determined by the structural characteristics of $\mathcal{E}$, and can be categorized as follows:
+where the extraction strategy $f_{\text{extract}}$ is determined by the structural characteristics of $\mathcal{E}$ and can be categorized as follows.
 Ego-centric
@@
 $V^{\prime}=\{v\in V\mid\text{dist}_{G}(v,v_{i})\leq k\}$ (5)
-, where $\text{dist}_{G}(u,v)$ denotes the shortest path distance between nodes $u$ and $v$ in $G$. The edge set $E^{\prime}$ is induced from $G$ by $V^{\prime}$.
+where $\text{dist}_{G}(u,v)$ denotes the shortest path distance between nodes $u$ and $v$ in $G$. The edge set $E^{\prime}$ is induced from $G$ by $V^{\prime}$.
 Multi-centric
@@
 The edge set $E^{\prime}$ is induced by $V^{\prime}$ from $G$.
-Appendix C Subgraph Pruning Algorithms
+Appendix C Subgraph Pruning and Visualization Strategy
 To maintain visual clarity and adhere to computational constraints, particularly for large graphs, the size of an extracted subgraph $G^{\prime}$ is bounded by a maximum node count, $N_{max}$.
-In instances where the initial extracted subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ exceeds this threshold, a principled pruning procedure is invoked to reduce its size while preserving the most topologically significant nodes.
-The specific pruning strategy is contingent upon the extraction method used, namely ego-centric or multi-centric, as detailed in Algorithm 1 and Algorithm 2, respectively.
+In instances where the initial extracted subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ exceeds this threshold, a principled pruning procedure is invoked.
+Furthermore, the remaining nodes are visualized using a task-driven strategy to ensure optimal perception by the VLM.
 C.1 Ego-centric Subgraph Pruning
@@
 This approach effectively reduces the subgraph’s complexity while guaranteeing that the core relational structure between the key entities remains intact.
+C.3 Task-Driven Subgraph Visualization Strategy
+The visual representation of the extracted subgraph $G^{\prime}$ plays a critical role in the VLM’s reasoning capability.
+We employ a task-driven visualization strategy where the graph layout algorithm is dynamically selected based on the task type $T$:
+  • Hierarchical Layouts: For sequential reasoning tasks such as Shortest Path or Cycle Detection, we utilize hierarchical layouts (e.g., Sugiyama algorithm).
+This layout emphasizes directionality and flow, enabling the VLM to visually trace paths layer-by-layer effectively.
+  • Force-Directed Layouts: For structural analysis tasks such as Community Detection or Clustering Coefficient, we employ force-directed algorithms (e.g., Fruchterman-Reingold).
+This approach naturally clusters densely connected nodes, making community structures visually distinct.
+Reproducibility and Clarity.
+To ensure strict reproducibility of the visual inputs, we fix the random seeds for all layout generation processes.
+Furthermore, to handle potential visual clutter in high-density subgraphs, we apply a visual anti-overlap mechanism.
+Following the pruning process, we dynamically adjust node repelling forces in the layout engine to ensure that no two nodes overlap and that edge crossings are minimized.
+This provides a clear “what-you-see-is-what-you-get” input for the Visual Graph Thoughts agent.
 Input : $G^{\prime}=(V^{\prime},E^{\prime})$: Initial $k$-hop neighborhood subgraph.
@@
 $V_{\text{cand}}\leftarrow V^{\prime}\setminus\{v_{i}\}$
-// Sort by priority: distance $\downarrow$, tier $\uparrow$, centrality $\uparrow$
-Sort $V_{\text{cand}}$ into list $L$ by key $(-d(v,v_{i}),t(v),c(v))$
+// Sort priority for removal: Farthest, Tier 3 (Peripheral), Lowest Centrality
+Sort $V_{\text{cand}}$ into list $L$ by key $(-d(v,v_{i}),-t(v),c(v))$
 $n_{\text{prune}}\leftarrow|V^{\prime}|-N_{\max}$
@@
 Model & Inference Configuration
-Model Path MODEL_PATH (Path to VLM)
 Tensor Parallel TENSOR_PARALLEL_SIZE 1
@@
 Our experiments are conducted using 8 NVIDIA H20 GPUs. The framework is implemented in Python 3.11 with CUDA Version 12.6, leveraging the vLLM library for efficient inference of the VLMs.
-Appendix F Implementation Details
+E.1 Experimental Settings and Hyperparameters
+Our experiments are conducted using 8 NVIDIA H20 GPUs.
+The framework is implemented in Python 3.11 with CUDA Version 12.6, leveraging the VLLM library for efficient inference of the VLMs.
 Key hyperparameters and framework configurations are detailed in Table 7.
+E.2 Process-Level DPO Training Details
+We fine-tuned $\mathcal{M}_{VRA}$ (instantiated with Gemma-3-12B and Qwen2.5-VL-7B) using the curated preference dataset $\mathcal{D}$.
+The DPO training was performed with a regularization parameter $\beta=0.1$.
+We utilized a learning rate of $5e^{-7}$ with a cosine decay scheduler and an effective batch size of 16 (achieved via gradient accumulation).
+To ensure memory efficiency during training, we employed LoRA (Low-Rank Adaptation) with rank $r=64$ and $\alpha=128$.
+The preference dataset $\mathcal{D}$ consists of 10,221 samples, specifically constructed to cover diverse reasoning failure modes.
+The distribution of error types in the rejected trajectories ($y_{l}$) is summarized in Table 9.
+Error Type Proportion Count
+Multimodal Errors 35% 3,577
+Textual and Logical Errors 65% 6,644
+Table 9: Distribution of error types in the rejected trajectories of the preference dataset.
+This balanced distribution ensures the model learns to robustly ground its reasoning in visual data while maintaining logical coherence.
 (a) Original graph
@@
 Figure 8: Visualizing Neighbor Retrieval for a target node A. The process identifies all nodes that are directly connected to node A.
+Appendix F Stability Analysis of DPO Training
+To further verify the robustness of the Visual Reasoning Agent ($\mathcal{M}_{VRA}$) within GraphVista, we conducted an in-depth analysis of the training stability of Process-level Direct Preference Optimization (DPO).
+This section presents detailed training metrics for Qwen2.5-VL-7B fine-tuned using DPO.
+F.1 Training Loss and Reward Curves
+Figure 9 illustrates the evolution of loss and implicit rewards over 7,500 training steps.
+(a) Training Loss
+(b) Implicit Rewards
+Figure 9: DPO training dynamics of Qwen2.5-VL-7B on the Grena preference dataset.
+The left panel shows rapid convergence of training loss; the right panel demonstrates that the model successfully widened the reward margin between Chosen and Rejected paths.
+Analysis of Training Dynamics
+The training process exhibited high stability without the divergence often observed in RL fine-tuning:
+  • Loss Convergence: The training loss (Figure 9(a)) displays a smooth downward trend, entering a stable convergence phase around 4,500 steps.
+This indicates that the model effectively learned the distribution of the preference data.
+  • Reward Margin Separation: As shown in Figure 9(b), the reward values for Chosen paths (indicated by the blue solid line) remain relatively stable, while the reward values for Rejected paths (indicated by the orange dashed line) decrease significantly with training steps.
+This indicates that the optimizer maximizes the preference margin primarily by suppressing negative samples (i.e., reasoning chains containing visual hallucinations or logical errors), aligning with the theoretical expectations of DPO.
+The observed stability is primarily attributed to our mild regularization strategy ($\beta=0.1$) and the high-quality process preference dataset derived from real topological structures.
 Appendix G Example of Visual Graph Thought Chains
@@
 Finally, Figure 8 depicts the fundamental operation of Neighbor Retrieval, where the model identifies and highlights all nodes directly adjacent to a given target.
-These examples demonstrate how Visual Graph Thoughts provide explicit, verifiable reasoning grounded in visual evidence.
-Appendix H Prompt and QA Templates
-This section provides prompt templates for evaluating the graph understanding capabilities of Vision-Language Models (VLMs) within the GraphVista framework.
-The templates are categorized into Simple Property Tasks and Complex Local Tasks, reflecting the task routing mechanism of the planning agent as described in Section 3.3 of the main paper.
-Each template includes an example prompt, the expected model behavior based on the GraphVista architecture, and the rationale behind the chosen modality.
-H.1 Simple Property Tasks: Text Modality with GraphRAG
-For simple property tasks, the planning agent routes the query to the text-modality branch.
-This branch utilizes GraphRAG-based retrieval on the Hierarchical GraphRAG Base ($\mathcal{K}$) to efficiently extract factual information without needing visual inference.
-H.2 Complex Local Tasks: Visual Modality with Visual Graph Thoughts
-For tasks requiring nuanced topological understanding or multi-step reasoning, the planning agent delegates the task to the visual modality branch.
-This module extracts a relevant subgraph, visualizes it, and employs the Visual Graph Thoughts agent to perform grounded, step-by-step multimodal reasoning.
+Task Categorization Examples 1.
+Text-Modality Tasks (Attribute, Statistic, Retrieval) Query: “What is the weight of the edge between Node 5 and 9?” Logic: Targets discrete attributes.
+Visual parsing is redundant; direct GraphRAG retrieval is optimal.
+2.
+Visual-Modality Tasks (Topological Reasoning) Query: “Identify a cycle of length 3 involving Node A.” Logic: Requires tracing paths.
+Visual inspection exploits spatial recognition, superior to textual adjacency processing.
+3.
+Collaborative Tasks (Global-Local Composition) Query: “Find the highest degree node and list its 2-hop neighbors.” Logic: Step 1 (Text): Global search for target node.
+Step 2 (Visual): Local subgraph reasoning.
+Figure 10: Categorization guidelines used by the Planning Agent to route tasks based on structural dependency profiles.
+Appendix H Task Parsing and Routing Mechanism
+This section delineates the decision-making logic of the Planning Agent, which serves as the semantic router within the GraphVista framework.
+Building upon the architectural overview, we detail the criteria and operational rationale for allocating tasks across Text, Visual, and Collaborative modalities.
+To balance computational efficiency with semantic flexibility, the Planning Agent employs a hierarchical two-stage routing protocol:
+  1. 1.
+    Deterministic Template Routing (Primary Strategy): The agent initially queries a library of predefined task templates (e.g., “What is the degree of node <<ID>>?”).
+Upon identifying an exact match, the task is deterministically typed and immediately routed to the designated modality, bypassing complex inference to minimize latency for standardized queries.
+  2. 2.
+    Semantic Inference Fallback (Secondary Strategy): In the absence of a template match, the agent resorts to semantic parsing.
+This mechanism analyzes the structural dependency of the query to infer user intent, specifically determining whether the solution necessitates topological traversal (Visual), attribute retrieval (Text), or a synergistic combination of both (Collaborative).
+Rationale for Text-Modality Assignment.
+The Text Modality is explicitly designated for tasks akin to database retrieval.
+This category encompasses: (1) Existence Verification, such as confirming the presence of nodes or edges; (2) Attribute Retrieval, which involves extracting specific properties (e.g., weights, labels) from the knowledge base $\mathcal{K}$; and (3) Global Aggregation, for operations like degree counting.
+For these tasks, direct retrieval offers computational superiority over visual processing, effectively mitigating latency and eliminating the noise associated with rendering dense graphs for simple factual extraction.
+Rationale for Visual-Modality Assignment.
+The Visual Modality is reserved for tasks necessitating "topological perception." Large Language Models (LLMs) processing linearized graph descriptions (e.g., adjacency lists) are prone to contextual attenuation ("lost-in-the-middle" phenomena) when tracing extended paths.
+Conversely, the visual modality projects the graph into a 2D layout where topological features—such as clusters, bridges, and cycles—become explicitly salient.
+This is critical for tasks like Shortest Path or Motif Detection, where the VLM can perceive connectivity patterns more effectively than through iterative textual deduction.
+Rationale for Collaborative-Modality Assignment.
+Complex graph reasoning often demands a hybrid "Search-then-Reason" paradigm.
+Collaborative tasks are characterized by a global search space (optimally handled by GraphRAG indexing) coupled with local structural verification (optimally handled by Visual Graph Thoughts).
+The Planning Agent decomposes such queries into sequential sub-goals.
+For instance, in diameter estimation, the Text branch first filters peripheral node candidates to reduce the search space, after which the Visual branch executes fine-grained pathfinding on the induced subgraph.
+H.1 Analysis of Task Classification Accuracy
+To evaluate the reliability of the routing mechanism, we quantify the task classification error rates across different VLM backbones.
+Table 10 presents the category misclassification rates for the Planning Agent.
+VLM Error Rate (%)
+GLM-4.1V-9B 0.1650
+InternVL3-9B 0.1593
+Qwen2.5-VL-7B 0.2482
+Gemma-3-12B 0.1697
+Qwen2.5-VL-7B (DPO) 0.2495
+Gemma-3-12B (DPO) 0.1713
+Table 10: Task Classification Error Rates across different VLM backbones. Models fine-tuned with DPO are denoted with (DPO).
+The results highlight two key observations regarding the routing stability:
+  • Impact of Model Scale: There is a clear correlation between model capacity and routing accuracy.
+Larger models, such as InternVL3-9B and GLM-4.1V-9B, consistently achieve lower error rates (approx.
+0.16%) compared to the 7B-parameter baselines (approx.
+0.25%).
+This suggests that the semantic nuance required for accurate intent classification benefits significantly from stronger foundational language understanding.
+  • Robustness to DPO Fine-tuning: The application of Direct Preference Optimization (DPO)—while crucial for enhancing the step-by-step visual reasoning detailed in the main methodology—does not degrade the semantic routing capability.
+The error rates for DPO-tuned models (e.g., Qwen2.5-VL-7B (DPO) at 0.2495) remain comparable to their base counterparts (0.2482).
+This indicates that our alignment strategy effectively isolates visual reasoning improvements without incurring an “alignment tax” on the model’s general instruction-following and planning abilities.
+System Prompt for Planning Agent Task Parsing and Routing Role Definition: You are an expert Graph Analysis Planning Agent.
+Your objective is to parse user queries regarding graph structures, extract key entities, classify the specific task type, and route the task to the optimal processing modality.
+Parsing Logic: 1.
+Template Verification: Initially, verify if the user input aligns with any predefined templates (e.g., Degree, Weight, Shortest Path).
+If a match is detected, output the standard JSON immediately.
+2.
+Semantic Fallback: Upon template mismatch, analyze the "structural dependency" of the request to determine the required modality (Text vs.
+Visual vs.
+Collaborative).
+Task Categorization Guidelines (for Fallback): Text-Modality Tasks: Tasks requiring direct retrieval of statistical data, node attributes, or global counts devoid of complex structural reasoning.
+Examples: Node Count, Node Degree, Edge Existence, Edge Weight.
+Visual-Modality Tasks: Tasks necessitating topological understanding, pathfinding, or pattern recognition within a local subgraph.
+Examples: Shortest Path, Cycle/Triangle Detection, Common Neighbors, Clique Detection.
+Modality-Collaborative Tasks: Tasks requiring global structural understanding necessitating decomposition into sequential text and visual sub-tasks.
+Examples: Graph Diameter, Critical Node Detection, Connectivity Analysis.
+Output Schema (JSON):
+  "task_type": "String (e.g., shortest_path, node_degree)",
+  "entities": ["List of key node IDs or null"],
+  "modality": "One of [Text, Visual, Collaborative]",
+  "reasoning": "Brief explanation of the modality choice",
+  "decomposition": ["List of sub-steps if modality is Collaborative, else null"]
+Few-Shot Demonstrations: User Input: "What is the degree of node 15?" Agent Output:
+  "task_type": "node_degree", "entities": ["15"], "modality": "Text",
+  "reasoning": "Template Match: ’node_degree’. Retrieves attribute via GraphRAG.",
+  "decomposition": null
+User Input: "Find the shortest path between node A and node B." Agent Output:
+  "task_type": "shortest_path", "entities": ["A", "B"], "modality": "Visual",
+  "reasoning": "Template Match: ’shortest_path’. Topology required.",
+  "decomposition": null
+User Input: "Analyze the network to find bottlenecks impacting flow." Agent Output:
+  "task_type": "bottleneck_detection", "entities": [], "modality": "Collaborative",
+  "reasoning": "Fallback: Unmatched template. Semantic parsing indicates global
+  search followed by local verification.",
+  "decomposition": [
+      "Identify high-betweenness candidates (Text)",
+      "Analyze local connectivity for flow restriction (Visual)"
+Current Request: <USER_QUERY>
+Figure 11: The system prompt utilized by the Planning Agent. The mechanism prioritizes template matching for efficiency, falling back to semantic parsing for complex or unstructured queries.
+Appendix I Prompt and QA Templates
+This section delineates the standardized prompt templates and Question-Answering (QA) protocols utilized to rigorously evaluate the graph structure understanding capabilities of Vision-Language Models (VLMs) within the proposed GraphVista framework.
+To ensure a comprehensive assessment of both retrieval accuracy and reasoning depth, we categorize these templates into Text-Modality Tasks and Visual-Modality Tasks.
+This categorization strictly adheres to the task routing logic executed by the Planning Agent, as detailed in Section 3.3 of the main paper.
+I.1 Text-Modality Tasks: Retrieval via GraphRAG
+The Text-Modality category encompasses tasks necessitating precise information retrieval and statistical aggregation rather than abstract topological reasoning.
+Upon classifying a query as text-dependent, the Planning Agent routes the task to the text-modality branch.
+This branch leverages the Hierarchical GraphRAG Base ($\mathcal{K}$) to execute targeted retrieval operations.
+By accessing the structured textual descriptions stored within $\mathcal{K}$ (e.g., node attributes, edge weights, and local neighborhoods), the model extracts factual information directly, thereby bypassing the potential ambiguity associated with visual inference for non-spatial queries.
+I.2 Visual-Modality Tasks: Reasoning with Visual Graph Thoughts
+This category targets tasks demanding intricate topological perception and multi-step logical inference, such as pathfinding or cycle detection.
+For such queries, the Planning Agent delegates execution to the visual-modality branch.
+This module orchestrates a pipeline consisting of: (1) dynamic subgraph extraction from $\mathcal{K}$ to isolate the region of interest; (2) task-driven visualization to render the subgraph into a high-resolution image; and (3) the deployment of the Visual Graph Thoughts Agent.
+This agent employs a chain-of-thought mechanism grounded in the generated visual evidence to perform iterative, step-by-step reasoning, ensuring that the final answer is topologically consistent.
 Task 1: Node Existence and Properties Prompt Template: "Determine if node <node_id> exists in the graph.
-If so, what is its degree?" GraphVista Behavior: The planning agent categorizes this as a simple property task.
+If so, what is its degree?" GraphVista Behavior: The planning agent categorizes this as a text-modality task.
 The model performs a targeted retrieval query against the GraphRAG Base $\mathcal{K}$ to find the entry for <node_id>.
@@
 Task 2: Edge Existence and Weight Prompt Template: "Is there a direct edge between node <node_id_1> and node <node_id_2>?
-If yes, provide its weight." GraphVista Behavior: The query is identified as a simple property task.
+If yes, provide its weight." GraphVista Behavior: The query is identified as a text-modality task.
 The system retrieves the adjacency information for <node_id_1> from $\mathcal{K}$.
@@
 The textual modality avoids the potential for visual ambiguity (e.g., overlapping edges or nodes) that could occur in a complex graph visualization.
-Task 3: Shortest Path Identification Prompt Template: "Find and list the sequence of nodes that form the shortest path between node <node_id_1> and node <node_id_2>." GraphVista Behavior: The planning agent categorizes this as a complex local task.
+Task 3: Shortest Path Identification Prompt Template: "Find and list the sequence of nodes that form the shortest path between node <node_id_1> and node <node_id_2>." GraphVista Behavior: The planning agent categorizes this as a visual-modality task.
 The Visual Graph Thoughts Agent extracts a multi-centric subgraph centered around the shortest path candidates between <node_id_1> and <node_id_2>.
@@
 Visual inspection allows the VLM to directly perceive the connectivity pattern, whereas a text-based approach would require inefficient and error-prone traversal of adjacency lists to check for path closures.
+Task 5: Graph Diameter Calculation (Modality-Collaborative) Prompt Template: "Calculate the diameter of the graph.
+First identify potential peripheral nodes, and then determine the maximum shortest path distance between them." GraphVista Behavior: The Planning Agent identifies this as a collaborative task and decomposes it into two sequential sub-tasks.
+Phase 1 (Text Modality): The agent queries the GraphRAG Base $\mathcal{K}$ to identify a set of "peripheral nodes" (typically defined as nodes with low centrality scores in Tier 3).
+This filters the search space to the most likely candidates for the diameter endpoints.
+Phase 2 (Visual Modality): The system extracts multi-centric subgraphs connecting these candidate pairs.
+The Visual Graph Thoughts agent then visually traces the shortest paths between them to calculate their distances.
+Finally, the system aggregates these results and returns the maximum distance found as the graph diameter.
+Rationale: Calculating diameter requires both global search and local pathfinding.
+The text modality is efficient for globally filtering candidate nodes based on statistical properties (centrality), avoiding the need to process the entire graph visually.
+Conversely, the visual modality excels at the subsequent topological reasoning required to accurately count path lengths between specific pairs, which is cumbersome for text-based reasoning.